### v2

# AgentA/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

Abstract.
A/B testing experiment is a widely adopted method for evaluating UI/UX design decisions in modern web applications. Yet, traditional A/B testing remains constrained by its dependence on the large-scale and live traffic of human participants, and the long time of waiting for the testing result. Through formative interviews with six experienced industry practitioners, we identified critical bottlenecks in current A/B testing workflows. In response, we present AgentA/B, a novel system that leverages Large Language Model-based autonomous agents (LLM Agents) to automatically simulate user interaction behaviors with real webpages. AgentA/B enables scalable deployment of LLM agents with diverse personas, each capable of navigating the dynamic webpage and interactively executing multi-step interactions like search, clicking, filtering, and purchasing. In a demonstrative controlled experiment, we employ AgentA/B to simulate a between-subject A/B testing with $1,000$ LLM agents Amazon.com, and compare agent behaviors with real human shopping behaviors at a scale. Our findings suggest AgentA/B can emulate human-like behavior patterns.
large language models, LLM agents, user experience, user interface, A/B testing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human-centered computing HCI design and evaluation methods
Figure 1. Architecture of AgentA/B system. The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) perform agent-web interaction with the assigned web environment, and 3) conduct post-testing analysis, and return the results to the end users.
[…]
A flowchart diagram illustrates the A/B testing workflow enabled by UXAgent. On the left, a system user (e.g., UX researcher or product manager) provides two types of input: LLM Agent specifications (e.g., number of agents, demographic distributions, attributes) and A/B Testing configuration (e.g., feature designs, agent intentions, metrics). These inputs lead to two parallel processes: LLM Agent Generation (Step 1) and Testing Condition Preparation (Step 2). The generated agents are then assigned to either a control or treatment group (Step 3), where they interact with their designated web environments. The control group uses the baseline interface, while the treatment group uses a modified (new) design. Simulation results are collected (Step 4) and passed to the Post-Testing Analysis module (Step 5), which returns A/B Testing Results (e.g., session outcomes, average actions, failure rates) back to the user. The visual includes labeled arrows, icons representing agents and interfaces, and text boxes describing each component.
1. Introduction
In the fast-paced online economy, user-facing web design decisions are central to user engagement and satisfaction. A/B testing–also known as online controlled experimentation–has become a cornerstone methodology for rapidly evaluating user interface designs and new system features (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015). A/B testing involves comparing two or more variants of a webpage (could be slight design changes or entirely new functionality) with real users randomly routed to the controlled group and the treatment (or experiment) group to evaluate the impact of design decisions on user engagements (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021). Major technology companies, including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), build their own A/B testing infrastructure and conduct hundreds of A/B tests to optimize their products and marketing strategies.
Despite being widely adopted, traditional A/B testing is inherently constrained by a number of participant-related challenges. First, online A/b testing requires a large sample size of users to achieve statistically significant results (often hundreds of thousands), but live user traffic is often expensive and resource-intensive, which may be especially problematic for niche markets or newer websites with low traffic (Kohavi et al., 2013; Xu et al., 2015). Second, user feedback cycles are inherently slow because a typical A/B test involves complex procedures such as experiment design, feature implementation, traffic allocation, result monitoring, and post-testing analysis (Fabijan et al., 2018), among which the result monitoring step alone may take weeks to complete. Third, the complexity of traffic coordination, engineering effort, and other administrative configurations means that the design teams can rarely test more than a handful of ideas throughout the design life-cycle. In practice, this means that many promising design variants never get tested due to operational constraints.
Recent advances in Large Language Model powered agents (LLM agents) offer a promising avenue to mitigate these limitations (Chen et al., 2025). LLM agents have demonstrated remarkable capabilities to mimic human decision-making, generate plausible behavior sequences, and reason across a range of domains, including healthcare (Li et al., 2024), software engineering (Luo et al., 2023), social science (Park et al., 2023; Hua et al., 2023), and autonomous agents (Wu et al., 2023; Yao et al., 2023). In particular, researchers have started to deploy LLM agents in web environments for tasks such as shopping (Yao et al., 2022), service booking (Zhou et al., 2023), and multimodal search (Koh et al., 2024). However, the vast majority of these systems operate in a single-agent session or work only on sandbox web environments, which often fail to capture the dynamic and diverse nature of the large-scale user interaction sessions on real-world websites.
In the context of A/B testing, we hypothesize LLM agents can help evaluate multiple design variants of web designs by role-playing real users’ interactions with an interactive web environment. This possibility could enable website designers and product managers to rapidly and risk-freely piloting many design candidates without worrying about limited user traffic. Our position is that LLM agents should not replace real user testings in a feature design project lifecycle; instead, it could complement the existing but limited piloting sessions for design optimization.
Our project began with a formative study with six experienced professionals who conduct A/B testing in their day-to-day workflow and extracted a current A/B testing workflow (Fig 2) and their challenges: 1) the lack of lightweight prototyping tools for early hypothesis testing, 2) the scarcity and contestation of user traffic as an experimental resource, and 3) the long turnaround times for user feedback that discourage design exploration. Based on this, we then developed AgentA/B, an LLM agent-based end-to-end A/B testing system, with the following four functions (Fig. 1): 1) LLM agent persona generation, 2) A/B testing condition preparation, 3) iterative and automated agent-web interaction, and 3) post-testing data analysis. Our system can plug and play with any existing LLM agent systems (e.g., Claude Computer-use agent or ReAct (Yao et al., 2023)) to generate hundreds of thousands of virtual users, each with a generated user persona, and these virtual users can interactively observe and operate on real webpages. For each agent-web interaction session (Fig. 3), our system has a web environment parsing module to preprocess dynamic web pages into structured, semantically meaningful JSON observations for LLM agents to consume, and has an action execution module to translate the LLM agent generated next action (e.g., “solar filter for telescope”) into a web operation (“click on solar filter” and “click first product”). The system monitors the LLM-agent reasoning logic and the agent behavior trajectories as a virtual user interaction session and automatically generates post-testing analysis results for users to analyze (e.g., t-test of purchased price in the two groups).
We use Amazon.com as a case study for automated A/B testing, where we generated 100,000 virtual customer personas and randomly selected $1,000$ to generate an LLM agent-based virtual customer using each of the personas. The design to be evaluated is whether the webpage shows all filter options in the left panel (control) or only shows a reduced list of options (treatment). The results show that LLM agents in the treatment group conducted more purchase and filter actions compared to those in the control group, which aligns with the direction of effects observed in the parallel human study. We also compared the behavior of LLM agents with aggregated results from 1 million human users and found that LLM agents were more goal-directed in their interactions, resulting in a shorter action length.
In summary, this paper implements AgentA/B, a system that deploys LLM-based agents into A/B testing conditions to interact with live websites to simulate realistic user interactions. Our case study of a 1,000 agent between-subject A/B testing on Amazon.com shows promising results of using the AgentA/B to conduct an automated and scalable Web A/B testing.
2. Related Work
We situate our work at the intersection of A/B testing, user behavior modeling, autonomous agents, and HCI research on interactive prototyping and simulation. We review four relevant threads of work: (1) limitations of traditional A/B testing practices, (2) advances in automated and optimized experimentation workflows, (3) LLM-based agent simulation across domains, and (4) interactive agents in web environments.
2.1. Limitations of Traditional A/B Testing in Practice
A/B testing has become a foundational methodology of comparing two versions of interfaces and functionalities of a webpage, app, or other digital assets to enable data-driven decisions about design strategies based on user behavior differences (Kohavi et al., 2009, 2007). The application of A/B testing spans both industry and academia, where organizations (e.g., major companies like Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017)) routinely rely on A/B testing to streamline the design, development, and deployment of products.
However, numerous studies in HCI communities have identified structural limitations in this widely adopted methodology. For instance, Fabijan et al. (2017) conducted an in-depth study across software-intensive organizations and found that A/B testing introduces slow iteration loops, costly feature development, and high failure rates—particularly when hypothesis formulation is weak. These findings were reinforced in a broader empirical survey by Fabijan et al. (2018), who emphasized the lack of actionable early-stage insights during the experimentation process. These works collectively underscore the HCI community’s recognition that while A/B testing provides rigor and promising benefits, it falls short in flexibility, speed, and insight generation.
Figure 2. The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites.
[…]
A timeline diagram illustrates the standard A/B testing workflow and highlights three key challenges. The timeline includes seven sequential phases: Feature Ideation (new UI layout or functionalities), A/B Testing Experiment Design (testing configuration and implementation), Data Monitoring (collect user interactions and rationales), Feature Deployment (final design decisions), Post-Testing Analysis (behavior patterns, failure cases, outcomes), Traffic Allocation (real user traffic toward variants), and Team Alignment (decision maker approval or rejection). The A/B Testing Period is marked across the Data Monitoring and Traffic Allocation stages. Three challenges are highlighted in red: (1) Scalable recruitment difficulties, (2) Limited testing opportunities, and (3) Slow user feedback loop. Each stage is connected along a central purple timeline.
2.2. Automated Experimentation and Interface Evaluation Tools
Given the aforementioned limitations, the HCI community has grown significant interest in systems that accelerate interface experimentation (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022; Swearngin et al., 2012). For example, Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) allowed designers to rapidly prototype for physical interaction such as touch interfaces. Fuse (Kuznetsov et al., 2022) enabled the rapid creation of context-aware UI behaviors from demonstrations.
On the experimentation side, Gilotte et al. (2018) proposed techniques for offline A/B testing, using logged user interaction data to estimate counterfactual outcomes for unobserved experimental variants. This reduces the need for live deployments but depends heavily on rich user logs and pre-existing infrastructure. Tamburrelli and Margara (2014) reframed A/B testing as a search problem within a design space, suggesting that evolutionary algorithms could be used to automate variant generation. These systems aim to reduce human effort in running and prioritizing experiments, yet still rely on historical data or user deployment. In contrast, our work explores agent-driven simulation as a complementary mechanism to automatically evaluate designs without live user data.
2.3. User Behavior Simulation: From Cognitive Models to LLM Agents
A longstanding tradition in HCI focuses on simulating user behavior through cognitive models such as GOMS and ACT-R (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996; Gray and Altmann, 2001). Building upon these traditional approaches, a number of works have leveraged inverse reinforcement learning (IRL) to infer user intentions from observed behaviors. Ziebart et al. (2009) introduce a probabilistic approach to IRL, enabling the modeling of complex user strategies in dynamic environments. However, these models are labor-intensive, require domain expertise, and do not generalize easily to open-ended tasks like online shopping. More recent work uses data-driven models to simulate user behavior in online environments. For example, Paranjape et al. (2016) used server logs to reconstruct navigation paths and optimize hyperlink structures. Oppenlaender et al. (2020) introduced CrowdUI, which used inverse reinforcement learning to infer task strategies from real user traces.
With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human-like behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024). Park et al. (2023) built a generative society of agents simulating daily social behavior in a simulated town. Horton (2023) evaluated whether LLMs could replicate outcomes from real behavioral experiments on prosociality. Furthermore, Lu et al. (2025a) demonstrates that fine-tuning large language models (LLMs) on real-world behavioral data significantly enhances their ability to generate accurate user actions. These systems suggest that LLMs can produce high-fidelity behavioral simulations across diverse tasks. Our work extends this direction into the visually rich, dynamic domain of live web interfaces, focusing on shopping behavior and automated A/B testing experiments.
2.4. LLM Agents in Web Environments
A growing body of work has investigated how autonomous agents can operate within web-based environments, driven by advancements in both machine learning and HCI communities. For instance, WebShop (Yao et al., 2022) introduced a shopping benchmark with templated webpages for studying goal-directed navigation. WebArena (Zhou et al., 2023) extended this to multi-domain service tasks and became a standard evaluation for web agent development. VisualWebArena (Koh et al., 2024) incorporated vision-language models to parse visual cues in interface design. Similarly, WebVoyager (He et al., 2024) and WILBUR (Lutz et al., 2024) emphasized task generalization and adaptive memory in agents interacting with open-ended web content.
However, all these systems operate within simulated environments, which, while useful for reproducibility, abstract away many complexities of the real webpage, such as dynamic loading, layout noise, unexpected modals, and inconsistent structure. Furthermore, the focus of existing work is primarily on task completion, not on evaluating or comparing the user experience across different design configurations. This gap reveals a growing need for systems that can adapt to the diversity and complexity of real-world web environments in which designers often have to balance trade-offs between design constraints and usability. Addressing this gap requires new system architectures and evaluation frameworks that can support interaction in the wild and enable comparative analysis of design decisions in terms of agent efficiency, robustness, and user-centered outcomes.
Participant Gender Job Title Team
P1 F Project Manager PM Team
P2 M Software Development Manager Engineer Team
P3 F Product Manager PM Team
P4 M Product Manager PM Team
P5 M Product Manager PM Team
P6 M Machine Learning Researcher Science Team
Table 1. Demographics of formative study participants.
3. Formative Study: Understanding Challenges in A/B Testing Workflows
3.1. Formative Study Method
To inform the design of the LLM agent-based A/B testing system, we conducted a formative study with six professionals experienced in designing and running A/B tests in industry settings. Through semi-structured interviews, we asked participants to recount a recent A/B test: how it began and ended, the stages involved, technologies used, and challenges encountered. We also probed how they addressed issues and why those strategies were chosen. These interviews helped us understand current practices, challenges, and user needs, which later informed our design decisions for the system.
We used a snowball sampling strategy (Biernacki and Waldorf, 1981) to recruit participants who currently work in the e-commerce industry and have experience designing and conducting A/B tests. We started with our friends, colleagues, and connections who fit the selection criteria and then asked them to refer their connection to participate in our study. All interviewees were located in the United States; A total of 6 participants (four product managers, one software development manager, and one machine learning researcher) were recruited to participate in the interview study (see Table 1 for details).
Interviews were conducted remotely, with audio recorded upon consent. Sessions lasted 45–62 minutes. We transcribed and de-identified all recordings before conducting a grounded theory analysis (Muller and Kogan, 2012). Two co-authors independently performed open coding on practices, tools, and challenges. Through collaborative discussion, we developed a shared codebook and organized codes into categories (e.g., experiment design, system bottlenecks), which were then synthesized into broader themes such as the A/B testing lifecycle and opportunities for system support. We finally applied the developed set of codes to the whole corpus of transcripts.
3.2. Formative Study Findings
In this section, we first report the overall workflow of an A/B testing project lifecycle (Fig.2), followed by the challenges and coping strategies reported by industrial practitioners. We conclude this section with design requirements for an LLM-Agent based simulation system for A/B testing.
A/B Testing Project Lifecycle
The A/B testing project life cycle, as reported in the interview data, unfolds across seven interdependent stages (Fig.2). It begins with new feature ideation, where ideas originate from individuals or teams and can range from minor UI changes to novel functionalities. Once proposed, the idea enters a team alignment and buy-in stage, involving collaborative discussion and refinement, culminating in approval from higher-level decision-makers. With alignment secured, the A/B testing experiment design stage follows, where critical aspects such as user segmentation, control/treatment conditions, success metrics, and evaluation criteria are meticulously planned—this stage precedes any development to ensure methodological rigor. Next, the feature deployment and iteration stage involves cross-functional collaboration to build and refine the feature, often demanding the most time and resources. Upon completion, the A/B test experiment launch integrates the feature into a live A/B testing environment where user exposure is randomized and monitored. After the test concludes, post-experiment data analysis is conducted with data scientists to assess whether success criteria were met and to understand the underlying reasons. Finally, the feature decision is made based on the outcome: successful features are launched, while unsuccessful ones are shelved.
Participants reported that the whole process can take anywhere from 3 months to a year. As aforementioned, the A/B test process is definitely a team sport, where different job roles need to collaborate with each other to design and develop the feature, and to design, launch, and analyze the A/B test experiment.
Challenges and Coping Strategies
Interviewees described several challenges in the current A/B testing workflow. One major issue is the high cost of new feature development, which often requires multiple engineers working for months to produce a functional version. During this lengthy process, it’s difficult to gather early user feedback to iterate on the design. The only formal feedback comes through the A/B test experiment, which may be too late. Some feature owners use colleagues as alpha testers, but this feedback is typically biased. Thus, interviewees expressed a need for lightweight prototyping and early-stage user testing methods that can offer feedback before committing to a formal A/B test.
“We had features … required changes across all [UI] stacks and beyond in other services, and it took one year and a half [to develop]…” (P2, Software Development Manager)
Another common challenge is competing for user traffic. When multiple teams want to test features that affect similar UI components, their experiments must be serialized, as running them in parallel would interfere with results. Currently, prioritization happens late in the process after development and just before the test launch, which could significantly delay the following steps. Interviewees called for a more data-driven way to manage and prioritize experiment launches.
“The biggest pain point, I think, is ‘traffic lanes’ … So experiments look at the same area; you cannot run them in parallel, so we have to split the [user] traffic. But we don’t have an automated way to split the traffic properly… ” (P2, Software Development Manager)
Some teams reported that up to one-third of experiments fail to meet their predefined success criteria. This high failure rate is frustrating because A/B testing in its current form leads to a one-shot decision: if a feature misses even one metric, the opportunity is lost. Interviewees expressed interest in predictive tools that could help forecast experiment outcomes so that they can revise features or experimental designs before the launch.
“The hypothesis behind [a failed feature design] that is it’s easier to access [the UI element when put it at the bottom] … and to match our competitors. [A/B test experiment shows] customer hates it, [because] they can no longer find the UI element that they used to and they quit. We saw the abandonment rate go through the roof [in the A/B test experiment].” (P3, Product Manager)
In summary, while A/B testing is widely adopted and supported by infrastructure such as user-routing platforms, interviewees identified several shortcomings: the high development cost, the lack of user traffic prioritization, and frequent experiment failure. They called for innovations that support early-stage prototyping, internal evaluation without real user exposure, and predictive assessment of test outcomes before the actual A/B testing period starts.
4. AgentA/B: Automated A/B Testing on the Web with LLM Agents
Based on feedback from the formative study, we designed and developed AgentA/B a modular system that enables automated A/B testing on live websites with LLM agents to simulate realistic and diverse user interactions. Traditional A/B testing is bottle-necked by a number of long-standing limitations, including the expensive requisites of large sample sizes and user traffic for significant results, the complicated and prolonged procedures for experiment preparation, execution, and analysis, as well as the limited resources that can only support experiments with very few designs/features. However, our system takes advantage of LLM agents for low-cost, rapid feedback loop, and high-throughput evaluation of interface variants without relying on live user deployment. The system is designed with modularity and extensibility in mind, which can be accommodated in different web environments, target user populations, and LLM agent architectures.
In this section, we describe the AgentA/B system architecture in detail, including the end-to-end A/B testing pipeline, the interaction loop of LLM agent and the web environment, and the implementation techniques that support robust and high-fidelity simulation possible on live web platforms.
4.1. System Overview and Pipeline
AgentA/B is designed as an end-to-end simulation system for LLM agent-based A/B testing in live browser environments. Analogous to the preparation stage of real user A/B testing, the end users of our system (i.e., UX researchers or product managers) need to determine details of A/B testing designs and provide two web environment variants to be tested. After taking the user inputs, the system operates on four LLM-powered modules (Fig. 1): LLM agent generation, testing preparation, autonomous A/B testing simulation, and post-testing analysis.
AgentA/B system users begin by specifying the LLM agent specifications and A/B testing configurations. The agent specification defines the target user population, including the number of agents, demographic and behavioral diversity (e.g., age, education, tech literacy), and other persona attributes. These personas drive the agents’ planning and reasoning processes and introduce behavioral variability across the simulation.
In parallel, users configure the A/B test by defining initial user intentions, behavioral metrics to track, and the design features to evaluate. User intentions guide the agent’s interaction trajectory and termination condition (e.g., “find a discounted Bluetooth speaker under \$30”). The design features (such as layout or interaction flow) are implemented in fully functional web environments. In our e-commerce scenario, the tested design feature is the layout of the filter panel on the website.
Once these inputs are provided by end users, the LLM Agent Generation Module queries the backend LLM to generate the specified number of LLM agents with diverse persona descriptions and intentions. The query explicitly instructs the LLM to ensure the generated LLM agent persona and intentions must follow the user-provided agent specifications. In our experiment in Section 5, we leveraged this module to generate 100,000 agents. We provide a sample persona generated by our system in Appendix A. After generating the LLM agents as the candidate A/B testing participants, the Testing Preparation Module performs the agent traffic allocation by splitting the agents into control (without new feature) and treatment (with new feature) groups, and each group is assigned to interact with the corresponding web environments. The statistics of agent characteristic distributions will be calculated within each group to ensure that the distribution of the LLM agents is relatively balanced; if the statistics are not balanced, the Testing Preparation Module will re-execute agent traffic allocation until satisfactory.
These web environments for both groups need to be launched using independent browser instances controlled through ChromeDriver (for web environment parsing) and Selenium WebDriver (for automated interaction execution) integration. During the automatic interaction with the web environment, each LLM agent begins to interact with the webpage using an autonomous action prediction loop, which is shown in Fig. 3. Each loop involves perceiving the current webpage state, interpreting the action space, predicting the next action, and executing that action in the browser. The details of this core automated agent-environment interaction process are explained below in Section 4.2. During the process, each step of the interaction is recorded, and the system monitors the overall session progression until termination, such that the agent accomplishes the intended goal or encounters failure cases.
After all agents complete their interactions with the assigned web environments, AgentA/B transitions to the result analysis stage. The Post-Testing Analysis Module is responsible for aggregating, interpreting, and presenting agent behaviors in a form that supports A/B-style experimental comparison. The output of this module serves as the primary feedback surface for the system user. Each agent session produces a fine-grained action trace that includes the full sequence of interactions, timestamps, webpage states, executed actions, intermediate rationales (when available), and final outcomes. These logs are collected asynchronously during the simulation and stored in a structured format. Upon session termination, the analysis module aggregates these records across both the control and treatment groups to extract comparative metrics and visualize key behavioral dynamics.
The post-test analysis module outputs summary statistics such as actions per session, session duration (in steps and time), and purchase completion rate. Researchers can also examine detailed behaviors (e.g., search or click filter usage) and compare them across A/B condition variants. The system supports stratified analysis by agent demographics or personas to identify subgroup differences. It automatically computes effect sizes (e.g., absolute/relative changes in conversion-like behaviors) to aid interpretation. For instance, when testing redesigned filters, the system can reveal whether agents refined searches more, completed tasks faster, and purchased more, which offers early insights on usability and adoption risks before live deployment.
Finally, AgentA/B maintains compatibility with external analytics pipelines. The result logs (in .JSON and .XSL) are exported in a format compatible with common data science tools, enabling downstream statistical modeling, significance testing, or integration with traditional A/B testing dashboards. Users can use AgentA/B to analyze thousands of simulated interaction session results into actionable insights for interface evaluation.
This pipeline allows AgentA/B to support fully automated LLM-agent-web interactions at scale under various environment configurations. A single experimental run can involve hundreds or thousands of sessions distributed across different personas and design conditions, all executed without any human intervention.
Figure 3. One action prediction iteration of the automated web testing in AgentA/B with an LLM agent. (1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session. In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces. (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration.
[…]
A system diagram shows how a language model-based agent interacts with a web environment. The process begins with a web interface (Control/Treatment) feeding into an Environment Parsing Module, which extracts the webpage as a JSON structure and defines an action space (e.g., search, filter, purchase, product_click, terminate). This information, along with a predefined persona and intention, is sent to the LLM Agent (Action Prediction), which outputs the next action (e.g., search “solar filter for telescope”). This action is executed through the Action Execution Module, which sends it back to the Web Environment, completing the interaction loop.
4.2. Agent-Environment Interaction Architecture
At the core of AgentA/B is an iterative mechanism where each LLM agent continuously interacts with a real web environment by dynamically updating its understanding of the environment and adjusting its actions accordingly. This architecture consists of three tightly integrated components: the Environment Parsing Module, the LLM agent, and the Action Execution Module. Together, these components enable robust operation in complex and dynamic web environments, for instance, our system evaluation in Section 5 is experimented on the Amazon [^1] http://www.Amazon.com platform.
Environment Parsing Module.
The interaction begins in a dedicated browser-based web environment session. Traditional approaches explored the extraction of webpages into screenshots or raw HTML representations. However, these approaches are impeded by extracting overly complicated information with a lot of unwanted raw webpage information. For instance, the raw HTML and DOM trees retain complicated hierarchical information of elements in a webpage, and the screenshots introduce irrelevant visual content and increase processing latency. In our work, the Environment Parsing Module in our system parses the web environment into structured observations with a JSON format that simplifies the structure of the website and stores only key information for the agent-web interactions. In particular, we use a ChromeDriver to execute a JavaScript processing script within the browser. This script selectively extracts targeted information directly from the raw HTML by extracting essential web elements.
For the e-commerce scenario, we specifically designed the script to extract web elements like product filters, titles, descriptions, and customer ratings based on their unique identifiers (IDs or classes). On the search result page, as shown in Fig. 4, for instance, we extract product details such as titles, names, ratings, reviews, and prices from the results section and gather filter options (e.g., Brand, Price, Delivery Day) from the left search panel. This approach is designed to remove irrelevant elements that are not relevant to the targeted design features and user interactions, such as advertisements, banners, or unrelated hyperlinks. The JSON file generated by the Environment Parsing Module provides a cleaner and more focused observation of the web environment to the LLM agent.
The Environment Parsing Module also identifies the current action space, which defines the set of allowable actions an agent can perform in the given context. These actions mimic the sequential steps a user would take while interacting with the website, such as typing keywords and clicking on items. Each action is structured in a format that the LLM agent can interpret and execute as part of its decision-making process.
The actions we define are represented in text format, which allows the agent to respond and perform tasks consistently. The key actions included are: (1) Search: The agent uses the search bar to find specific items or information. (2) Click Product: Select an item from the current webpage to view more detailed information about it. (3) Click Filter Option: Apply one of the available filters (e.g., price range, brand) to refine the search results. (4) Purchase: Complete the purchase of the selected item. (5) Stop: Indicate that the shopping session is complete and no further actions are required. This approach supports realistic simulation of shopping processes, providing a basis for evaluating LLM capabilities and examining human-like behavior in online shopping environments.
LLM Agent.
The structured webpage and action space are passed to the LLM agent, which also receives the agent’s persona and intention–the initial intention was generated by the LLM Agent Generation Module but could be dynamically updated within the LLM Agent. The intention specifies the agent’s current task (e.g., searching for a specific product, comparing alternatives, or making a budget-constrained purchase).
The LLM agent functions as a decision-making module that consumes the current state and outputs the next action to be taken. In particular, the LLM agent models the next-step decision-making problem as a form of language-based reasoning and planning task by mapping structured state observations into reasoning traces and action predictions. AgentA/B is not bound to a specific LLM agent implementation. Instead, our system treats the LLM agent as an exchangeable module that supports various types of LLM web agents (ReAct (Yao et al., 2023), FireClaw [^2] https://github.com/mendableai/firecrawl-mcp-server) with convenient “plug-and-play” APIs, analogous to the Model Context Protocol (MCP) proposed by Claude. In our experiment, we adopt the UXAgent framework (Lu et al., 2025b), a state-of-the-art LLM agent for web interactions, as a representative implementation due to its strong performance and support for multi-step planning and intermediate memory.
Figure 4. The two design variants of the left filter panel on Amazon.com for the A/B testing case study with AgentA/B. In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).
[…]
A side-by-side comparison of two Amazon search results pages for the query ”solar filter for telescope.” Both pages show similar products but have differences in their filter options on the left sidebar. The left image displays additional filter categories under ”Department” such as ”Telescope Refractors,” ”Safety Goggles & Glasses,” and ”Telescope Eyepieces,” which are not present in the right image. Both images include nested subcategories under ”Camera & Photo Filters” and identical product listings on the right side of the page. This comparison highlights subtle differences in the user interface that could affect user navigation or filtering behavior.
Action Execution Module.
The next action predicted by the LLM agent is translated into browser commands by the Action Execution Module. Actions are expressed in a structured format that can reference DOM elements or logical operations, such as Click_product(3), Click_filter_option(Brand: Sony), Search(”Wireless earbuds”), or Purchase. The execution module parses the action and performs the corresponding interaction on the live webpage. In some cases, the web execution is not guaranteed to succeed, as real-world pages are prone to dynamic content loading and modal interruptions. Therefore, the execution module incorporates built-in fault detection and recovery logic. If an action fails due to a missing selector or DOM mismatch, the system attempts fallback options such as retrying, scrolling into view, or re-parsing the page. Each successful or failed execution updates the environment and starts the next iteration of the loop.
This interaction loop continues until the agent reaches a terminal condition. Successful termination occurs when the agent completes its task—for example, by navigating to a purchase page or explicitly declaring task success. Failure conditions include endless loops, unreachable goals, or repeated in-executable actions. Sessions are also capped with time and action count thresholds to prevent infinite rollouts. Each completed session produces a full trace of interaction history, action rationales, page states, and final outcomes.
5. Case Study: A/B Testing with Reduced Filter Options on Amazon.com
To evaluate the effectiveness of AgentA/B, we conducted large-scale simulations on a live e-commerce platform.
5.1. Method: A/B Testing Scenario
We designed a simulated A/B testing scenario focused on the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce. The goal was to evaluate how variations in filter configurations influence user behavior.
{booktabs}
colspec=lccc Control Condition
Full Filter List
Human, N=1M Control Condition
Full Filter List
Agent, N=500 Treatment Condition
Reduced Filter List
Agent, N=500
Search 6.40 1.42 1.43
Click_product 6.96 1.87 2.09
Click_filter_option 0.33 0.58 0.60
Purchase 0.62 0.81 0.83
Stop - 0.19 0.17
Average actions 15.96 6.05 6.60
# of purchase * - 404 414
Average \$ spend - \$55.14 \$60.99
Table 2. Comparison of human customers’s actions per session in control condition (current design with full filter list), and the LLM agents as virtual customers in both control condition and treatment condition (new design with reduced filter list). The LLM agents in treatment condition purchased significantly more items than the ones in the control condition. $\chi^{2}(1)=5.51$, p-value ¡ 0.05
Our team has the privilege of altering the left filter panel interface (Fig. 4). In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the users’ search query.
Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions. Each agent was initialized with a persona profile and a shopping goal (e.g., “find a budget smart speaker under \$40 with strong customer reviews”). The allowed action space includes typical e-commerce flows: search, filter use, product examination, and cart actions. Persona generation followed the methodology in Chen et al. (2024) with LLM agent’s demographics (age, income, occupation), preferences, and shopping goals.
In the case study, our AgentA/B was implemented with the Claude 3.5 Sonnet model as the LLM backend to support agent generation, testing preparation, automated agent-web interaction, and post-testing analysis. The AgentA/B environment was executed on a distributed cluster of 16 high-memory compute nodes, with each node controlling a Selenium-driven Chrome instance running in headless mode. Each session was capped at 20 actions. Sessions ended either with task success (completion or no further predicted actions) or failure (e.g., looping behavior). We logged full action traces, metadata (e.g., duration, outcome), and agent rationales where applicable.
5.2. Finding: Alignment with Human Behavior
To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using AgentA/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions. Table 5.1 presents the aggregated actions and statistics from sessions with LLM agents and with human participants.
Human participants engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs. 6.05). They also clicked more products (6.96 vs. 1.87) and performed more searches (6.40 vs. 1.42). Despite these differences in interaction style, humans and agents showed similar purchase rates (0.62 vs. 0.81) and comparable use of filter options (0.33 vs. 0.58).
These findings highlight notable behavioral differences: while LLM agents adopt more goal-directed, structured interaction strategies, human users display broader exploratory behaviors. Nevertheless, the decision-making processes and general interaction patterns of the agents reasonably approximate human behavior for intention-driven tasks. This alignment supports the utility and validity of employing agent simulations for controlled UX evaluations, especially in scenarios where rapid interface assessment without live deployment is beneficial.
5.3. Findings: System Effectiveness Across Interface Variants
To assess whether AgentA/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations. As shown in Table 5.1, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
Agents in the treatment condition clicked on more products (2.09 vs. 1.87), and the sessions have a higher total number of actions (6.60 vs. 6.05, t(998)=1.08, p=0.28). Most notably, the LLM agents in the treatment condition click filter options more frequently than the ones in the control condition (0.60 vs. 0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability. Lastly, when looking at the final outcome of the sessions in the treatment condition and control condition, LLM agent simulated virtual customers purchased more products (414 v.s. 404, $\chi^{2}(1)=5.51$, p=0.03) , and spent more money (\$60.99 v.s. \$55.14, t(998)=0.39, p=0.69) than the virtual customers in the control condition. Most of these comparisons, except the “# of product purchase” are not statistically significant, which suggests to the A/B test owner that this filter interface design demands a higher number of participants size (N). Importantly, the LLM-agent A/B testing results are consistent with the direction of changes in the human A/B testing results (which can not reported due to privacy policy).
In summary, the A/B simulation conducted with AgentA/B reveals that LLM-based agents, when equipped with structured personas and embedded intentions, produce behaviorally rich and interface-sensitive interactions. These behaviors vary in predictable ways based on persona, align with human user baselines, and are capable of distinguishing between subtle interface design variants. These findings validate the system’s use as a behaviorally grounded, scalable alternative to conventional A/B testing, suitable for early-stage prototyping, pre-deployment design validation, and hypothesis-driven UX evaluation.
6. Discussion
Our design and evaluation of AgentA/B demonstrate the feasibility and utility of using LLM-based autonomous agents at scale for behavioral simulation in the life-cycle of the web interface designs. These findings echo a growing trend in Machine Learning and HCI, where large-scale LLM agent-based simulations are increasingly explored as tools for social simulations, economic experiments, and iterative design support (Park et al., 2023; Horton, 2023; Gao et al., 2024). In this section, we discuss the broader implications of our findings, system-level affordance, current limitations, and future research directions.
6.1. Accelerating Design Iteration Through Simulated Feedback Loops
One of the most immediate and impactful contributions of AgentA/B lies in its ability to expedite the feedback loop for A/B testing and design evaluation in web-based applications. Traditional A/B testing, despite remaining the gold standard for user-centered and evidence-based design validation approaches, suffers from well-documented challenges that are primarily related to the experiment participants. In particular, real user traffic is often limited and expensive, feedback is slow due to the long time span for experiment execution, and there are not enough resources to test all the design variants (Fabijan et al., 2017; Kohavi et al., 2013; Xu et al., 2015). These structural constraints make it difficult for teams to validate novel ideas, particularly during early prototyping stages, when feedback is most critical. Furthermore, many practical challenges in engineering time and testing infrastructure often restrict the number of A/B testings that can be executed.
Recent work has emphasized the importance of early-stage, actionable feedback in early design phases to support interface and feature design (Fabijan et al., 2018). Tools like Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) have emphasized the value of rapid prototyping. However, most of these tools focus on interface prototyping rather than enabling feedback that is grounded in realistic user behavior.
Our AgentA/B, on the contrary, fills this critical gap by introducing a new phase into the design life-cycle: agent-based pilot experiments before real user testing. Our system enables a rapid and automated feedback loop through two key technical capabilities: 1) automatically scaling up massive amounts of LLM agents with realistic and diverse personas, and 2) asking the agents to perform realistic behaviors by “role-playing” real humans. These simulations are executed within live browser environments, with LLM agents continuously perceiving structured page representations, reason through goals and interface states, and predicting plausible next actions. The predicted actions are automatically translated into executions in web environments to picture individual user journeys. As a result, AgentA/B users, such as UX designers, can obtain large-scale, fine-grained insights into user decision strategies with rationales, failure cases, and behavioral divergence across design variants. These insights can support end users in iterating the designs and revising the actual user A/B tests before any real human participants are recruited or any feature is deployed.
6.2. Inclusive Piloting and Risk-Free Testing for Underrepresented Populations
In addition to accelerating iteration, AgentA/B enables another critical capability: inclusive, risk-free piloting for user populations that are otherwise difficult to recruit or ethically sensitive to test. These populations (e.g., older adults) are often underrepresented in early testing phases due to ethical or logistical barriers, and a poorly designed application may even impose harm on the user participants.
AgentA/B allows designers to simulate interactions from such groups by configuring agent personas that reflect the demographic and behavioral characteristics of targeted populations. For example, agents can be instantiated to simulate limited digital literacy, slower decision-making, or specific accessibility needs to reflect the behaviors of older adults more realistically and believably. By using our AgentA/B to generate a targeted agent population at scale and, more importantly, to automatically and iteratively test different design variants, designers can not only better evaluate how these groups might experience proposed features in real user testing but also fix design flaws and avoid any potential risks.
This capability not only enhances safety and inclusion in the design process but also ensures that interface and feature designs are more robust and accessible. Our idea echoes the early exploration of scaling interface evaluation using crowd-based approximations of end-user behavior (Bernstein et al., 2011; Oppenlaender et al., 2020). Our findings suggest that even in early prototyping phases, agents can offer intent-driven behavior patterns that may disclose usability bottlenecks, task friction, or potential failures. This paradigm shift reframes LLM agent simulation not merely as a performance benchmark but as a core design methodology. Agent-based piloting complements traditional user testing with risk-free experimentation, broadened design coverage, and prompt usability feedback that aligns with realistic interaction patterns.
6.3. Behavioral Fidelity of LLM Agents
Our results show that LLM-based agents can generate realistic behavior trajectories that are aligned with various human shopping strategies. The alignment in task completion rates, refinement behaviors, and interaction flow with real human data reflects the growing consensus in HCI that advanced LLMs can perform believable and context-aware behaviors (Park et al., 2023; Horton, 2023; Lu et al., 2025a).
However, simulated behavior is not a complete substitute for human cognition. As prior work in cognitive modeling and HCI simulation has emphasized (Gray and Altmann, 2001; John and Kieras, 1996; Ziebart et al., 2009), real humans are influenced by complicated explicit and implicit factors, including emotion, fatigue, prior experience, and latent goals. Many of these factors are not accessible to the current LLM agents. For instance, in our evaluations, agent behavior tended to be more deterministic, focused, and efficient compared to human users, who often displayed broader exploration and casual browsing. The behavior trajectories, as well as reasoning generated by LLM agents, should not be interpreted as perfect replications of user behaviors but as structured approximations that emphasize user behavior coverage and consistency. Our idea is analogous to the use of cognitive and computational models in prior HCI work, such as GOMS and ACT-R, which were not designed to replicate human behaviors perfectly but to estimate the interaction trade-offs at scale (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996).
6.4. Toward Simulation-Supported Design Methodologies and Automated Design Optimization
While AgentA/B currently focuses on simulating A/B testing between interface design variants, our system architecture supports broader capabilities that move beyond this particular evaluation methodology. Our system laid the foundation for AI-based human behavior simulation at scale to support various design explorations, refinements, and evaluations. For instance, we can adapt our system to cognitive walkthrough and heuristic evaluation by designing the LLM agents to role-play different types of stakeholders (e.g., domain experts, HCI designers, technical developers, etc.) and specifying the agent’s intention to be aligned with the design methodologies correspondingly. This novel opportunity echoes a number of established works that underscore the development and adoption of automated tools for interface exploration and optimization promptly, such as Fuse (Kuznetsov et al., 2022), Apparition (Lasecki et al., 2015), and d.tools (Hartmann et al., 2006).
Further, our AgentA/B sheds light on the vision of “design mining” (Kumar et al., 2013), where large numbers of alternatives are evaluated computationally to identify the most promising solutions. AgentA/B supports this approach by serving as a simulation engine for a large pool of candidate variants and automating the design optimization through full interaction traces. For example, if $50$ layout variants of a filter panel were proposed, agent simulations could conduct large-scale simulations and rank the designs by search efficiency, user engagement, or failure rate across different user subgroups.
This extension of the system functions beyond as a testbed but as an interactive co-design partner to provide empirical evidence throughout the life-cycle of novel designs and functionalities. We view this direction as a natural evolution of the current system architecture and an exciting opportunity for future integration with generative design tools and adaptive interface frameworks.
6.5. Limitations and Future Directions
While AgentA/B establishes a strong foundation for the scalable and automated LLM agent-based A/B testing, we acknowledge two key limitations in the current scope of the system and inform ongoing development.
First, the robustness of our system and the behavioral fidelity of simulated agents remain bound by the reasoning and grounding capabilities of the underlying LLMs. Although recent LLMs such as Claude 3 and GPT-4o exhibit exceptional language understanding and planning capacity, they can still misinterpret complex or unconventional DOM structures, particularly in dynamically rendered web environments. Even state-of-the-art LLMs may fail to correctly parse web environments when encountering inconsistent element labeling, interactive latency, or unexpected modal interruptions (Zhou et al., 2023; Koh et al., 2024). While our system mitigates these issues through a novel implementation of leveraging JSON for structured environment parsing, LLM agent performance can still degrade when real-time content, along with the web environment structures, continuously updates and changes.
Second, agent behaviors in the current AgentA/B system are not able to comprehensively incorporate affective or meta-cognitive signals such as uncertainty, fatigue, or emotional response. These dimensions could be challenging for LLMs to model but are critical in shaping human cognitive behaviors and decision-making (Horton, 2023; Ziebart et al., 2009). LLM agents in our system tend to simulate goal-driven behavior with plausible cognitive structure but do not yet capture the full range of human variability or intent ambiguity.
Beyond these limitations, several technical and conceptual opportunities emerge for future work. One promising direction lies in expanding agent capabilities through multimodal information perception. For instance, integrating visual inputs (e.g., screenshots, spatial layout) with text-based content could allow agents to operate more robustly across richly designed or accessibility-diverse interfaces but still may introduce irrelevant information along the path. Recent progress in vision-language models and multimodal agents (Gupta and Kembhavi, 2023; Chezelles et al., 2024) suggests a pathway to generalizing agent behaviors beyond text-only interfaces.
Another important opportunity is collaborative or multi-agent simulation, especially for applications such as shared productivity tools, learning platforms, or social platforms. In these scenarios, agent interactions may involve negotiation, coordination, or conflict. Building on work like generative agents for social simulations (Park et al., 2023), future systems could simulate rich user ecosystems rather than independent task-solving to better reflect the collaborative work nature in real-world scenarios.
These directions demonstrate the extensibility of AgentA/B and its potential to evolve from a simulation platform into a broader foundation for behaviorally informed, user-model-driven design tools. As LLM-based interaction models continue to improve, we anticipate these agents will not only validate designs but also help co-create them.
7. Conclusion
In this paper, we presented AgentA/B, a system that enables large-scale, LLM agent-based simulation of A/B testing for web interfaces in real browser-based web environments. Our evaluation demonstrated that LLM agents exhibit realistic, goal-aligned behaviors, are sensitive to interface variations, and provide actionable feedback comparable to human users. By supporting rapid, risk-free behavioral piloting, AgentA/B introduces a new phase for agent-based piloting in the design life-cycle that complements traditional A/B testing and expands the scope of early-stage UX evaluation. We envision future extensions that further enhance agent fidelity, broaden domain coverage, and integrate simulation into intelligent design optimization workflows.
Appendix A Example of Persona Generated by AgentA/B
{minted}
[breaklines, breaksymbol=,breakanywhere]text Persona: Marcus
Background: Marcus is a 35-year-old freelance graphic designer living in Austin, Texas. After working for a decade in various creative agencies, he transitioned to freelancing to gain more control over his schedule and focus on passion projects, such as illustrating indie game assets and creating digital art for local musicians.
Demographics:
Age: 35
Gender: Male
Education: Bachelor’s degree in Visual Communication
Profession: Freelance Graphic Designer
Income: $70,000(variablebasedonprojects)\par FinancialSituation:Marcusearnsadecentlivingfromhisfreelancegigs,thoughhisincomecanfluctuate.He^{\prime}sfinanciallystablebutcautiousaboutbigexpenses.Hesetsasidepartofhisearningsfortravelandsoftwareupgrades,whichareessentialforhiswork.\par ShoppingHabits:Marcusenjoysdiscoveringuniqueornicheproducts,especiallytechgadgets,artsupplies,andstreetwear.Heprefersshoppingonlineforthevarietyandreadsreviewscarefully.He^{\prime}sbrand-loyalwhenitcomestotoolsherelieson,likehisdrawingtabletanddesignsoftware.Forclothing,he^{\prime}sdrawntobold,graphic-heavyitemsthatreflecthisartisticvibe.\par ProfessionalLife:Marcusworksfromahomestudiothatdoublesasacreativespace.Hecollaboratesremotelywithclientsfromdifferentindustries,jugglesmultipledeadlines,andoftenpullslatenights.Hefrequentlyupdateshisportfolioandmaintainsastrongsocialmediapresencetoattractnewclients.\par PersonalStyle:Marcushasanedgyandexpressivefashionsense.Hewearslarge-sizedclothingandgravitatestowarddarktoneswithpopsofneonorgraphicprints.Comfortisimportant,buthelikestomakeavisualstatementwithwhathewears.Hisgo-tooutfitisasofthoodiewithacustomdesign,blackjoggers,andhigh-topsneakers.\par\@add@PDF@RDFa@triples\par$

### v3

@@
 Abstract.
-A/B testing experiment is a widely adopted method for evaluating UI/UX design decisions in modern web applications.
-Yet, traditional A/B testing remains constrained by its dependence on the large-scale and live traffic of human participants, and the long time of waiting for the testing result.
-Through formative interviews with six experienced industry practitioners, we identified critical bottlenecks in current A/B testing workflows.
-In response, we present AgentA/B, a novel system that leverages Large Language Model-based autonomous agents (LLM Agents) to automatically simulate user interaction behaviors with real webpages.
-AgentA/B enables scalable deployment of LLM agents with diverse personas, each capable of navigating the dynamic webpage and interactively executing multi-step interactions like search, clicking, filtering, and purchasing.
-In a demonstrative controlled experiment, we employ AgentA/B to simulate a between-subject A/B testing with $1,000$ LLM agents Amazon.com, and compare agent behaviors with real human shopping behaviors at a scale.
-Our findings suggest AgentA/B can emulate human-like behavior patterns.
+A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long run times, and high operational cost.
+To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable and behavioral evidence before launch.
+In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases (414 vs.
+403), reproducing directional outcomes also observed in a parallel large-scale human experiment.
+Results further suggest that agent-based simulations can detect interface-sensitive behavioral differences and surface subgroup patterns while offering faster, lower-risk insights.
+We position Agent A/B as a complement to human testing, enabling earlier prototyping, pre-deployment validation, and hypothesis-driven UX evaluation.
 large language models, LLM agents, user experience, user interface, A/B testing
@@
 Architecture of Agent A/B system.
-The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) perform agent-web interaction with the assigned web environment, and 3) conduct post-testing analysis, and return the results to the end users.
-[…]
-A flowchart diagram illustrates the A/B testing workflow enabled by UXAgent.
-On the left, a system user (e.g., UX researcher or product manager) provides two types of input: LLM Agent specifications (e.g., number of agents, demographic distributions, attributes) and A/B Testing configuration (e.g., feature designs, agent intentions, metrics).
-These inputs lead to two parallel processes: LLM Agent Generation (Step 1) and Testing Condition Preparation (Step 2).
-The generated agents are then assigned to either a control or treatment group (Step 3), where they interact with their designated web environments.
-The control group uses the baseline interface, while the treatment group uses a modified (new) design.
-Simulation results are collected (Step 4) and passed to the Post-Testing Analysis module (Step 5), which returns A/B Testing Results (e.g., session outcomes, average actions, failure rates) back to the user.
-The visual includes labeled arrows, icons representing agents and interfaces, and text boxes describing each component.
+The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) prepare the testing conditions (e.g., traffic allocation), 3) perform agent-web interaction with the assigned web environment, 4) monitor and record agent behaviors as well as web interactions, and 5) conduct post-testing analysis so that the results will be returned to the end users.:
 1. Introduction
-In the fast-paced online economy, user-facing web design decisions are central to user engagement and satisfaction.
-A/B testing–also known as online controlled experimentation–has become a cornerstone methodology for rapidly evaluating user interface designs and new system features (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015).
-A/B testing involves comparing two or more variants of a webpage (could be slight design changes or entirely new functionality) with real users randomly routed to the controlled group and the treatment (or experiment) group to evaluate the impact of design decisions on user engagements (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021).
-Major technology companies, including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), build their own A/B testing infrastructure and conduct hundreds of A/B tests to optimize their products and marketing strategies.
-Despite being widely adopted, traditional A/B testing is inherently constrained by a number of participant-related challenges.
-First, online A/b testing requires a large sample size of users to achieve statistically significant results (often hundreds of thousands), but live user traffic is often expensive and resource-intensive, which may be especially problematic for niche markets or newer websites with low traffic (Kohavi et al., 2013; Xu et al., 2015).
-Second, user feedback cycles are inherently slow because a typical A/B test involves complex procedures such as experiment design, feature implementation, traffic allocation, result monitoring, and post-testing analysis (Fabijan et al., 2018), among which the result monitoring step alone may take weeks to complete.
-Third, the complexity of traffic coordination, engineering effort, and other administrative configurations means that the design teams can rarely test more than a handful of ideas throughout the design life-cycle.
-In practice, this means that many promising design variants never get tested due to operational constraints.
-Recent advances in Large Language Model powered agents (LLM agents) offer a promising avenue to mitigate these limitations (Chen et al., 2025).
-LLM agents have demonstrated remarkable capabilities to mimic human decision-making, generate plausible behavior sequences, and reason across a range of domains, including healthcare (Li et al., 2024), software engineering (Luo et al., 2023), social science (Park et al., 2023; Hua et al., 2023), and autonomous agents (Wu et al., 2023; Yao et al., 2023).
-In particular, researchers have started to deploy LLM agents in web environments for tasks such as shopping (Yao et al., 2022), service booking (Zhou et al., 2023), and multimodal search (Koh et al., 2024).
-However, the vast majority of these systems operate in a single-agent session or work only on sandbox web environments, which often fail to capture the dynamic and diverse nature of the large-scale user interaction sessions on real-world websites.
-In the context of A/B testing, we hypothesize LLM agents can help evaluate multiple design variants of web designs by role-playing real users’ interactions with an interactive web environment.
-This possibility could enable website designers and product managers to rapidly and risk-freely piloting many design candidates without worrying about limited user traffic.
-Our position is that LLM agents should not replace real user testings in a feature design project lifecycle; instead, it could complement the existing but limited piloting sessions for design optimization.
-Our project began with a formative study with six experienced professionals who conduct A/B testing in their day-to-day workflow and extracted a current A/B testing workflow (Fig 2) and their challenges: 1) the lack of lightweight prototyping tools for early hypothesis testing, 2) the scarcity and contestation of user traffic as an experimental resource, and 3) the long turnaround times for user feedback that discourage design exploration.
-Based on this, we then developed AgentA/B, an LLM agent-based end-to-end A/B testing system, with the following four functions (Fig.
-1): 1) LLM agent persona generation, 2) A/B testing condition preparation, 3) iterative and automated agent-web interaction, and 3) post-testing data analysis.
-Our system can plug and play with any existing LLM agent systems (e.g., Claude Computer-use agent or ReAct (Yao et al., 2023)) to generate hundreds of thousands of virtual users, each with a generated user persona, and these virtual users can interactively observe and operate on real webpages.
-For each agent-web interaction session (Fig.
-3), our system has a web environment parsing module to preprocess dynamic web pages into structured, semantically meaningful JSON observations for LLM agents to consume, and has an action execution module to translate the LLM agent generated next action (e.g., “solar filter for telescope”) into a web operation (“click on solar filter” and “click first product”).
-The system monitors the LLM-agent reasoning logic and the agent behavior trajectories as a virtual user interaction session and automatically generates post-testing analysis results for users to analyze (e.g., t-test of purchased price in the two groups).
-We use Amazon.com as a case study for automated A/B testing, where we generated 100,000 virtual customer personas and randomly selected $1,000$ to generate an LLM agent-based virtual customer using each of the personas.
-The design to be evaluated is whether the webpage shows all filter options in the left panel (control) or only shows a reduced list of options (treatment).
-The results show that LLM agents in the treatment group conducted more purchase and filter actions compared to those in the control group, which aligns with the direction of effects observed in the parallel human study.
-We also compared the behavior of LLM agents with aggregated results from 1 million human users and found that LLM agents were more goal-directed in their interactions, resulting in a shorter action length.
-In summary, this paper implements AgentA/B, a system that deploys LLM-based agents into A/B testing conditions to interact with live websites to simulate realistic user interactions.
-Our case study of a 1,000 agent between-subject A/B testing on Amazon.com shows promising results of using the AgentA/B to conduct an automated and scalable Web A/B testing.
+A/B testing—also known as online controlled experimentation—has become a cornerstone methodology for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015).
+By comparing alternative interface designs or feature functionalities with real users randomly assigned to control and treatment groups, A/B testing provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021).
+Major technology companies,including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), rely on hundreds of such tests each year to refine features, improve engagement, and guide strategic design choices.
+Despite its ubiquity, however, traditional A/B testing is often slowed by practical constraints, including heavy reliance on live user traffic, long turnaround times for statistically reliable results, and significant engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
+To better understand these challenges, we conducted a formative study with six industry practitioners who regularly run A/B tests in their day-to-day workflows.
+Their accounts highlight three recurring bottlenecks (Fig 2).
+First, there is a lack of lightweight methods for piloting design hypotheses before committing to large-scale traffic allocation.
+Second, user traffic itself is a scarce and contested resource, particularly for niche products or teams with competing priorities.
+Third, feedback cycles are slow: weeks may pass between experiment setup and actionable insights, discouraging exploration of alternative designs.
+Collectively, these constraints mean that many promising ideas are never tested, and early-stage prototyping rarely benefits from rigorous, behavioral-data grounded evaluation.
+Recent advances in large language model (LLM)–based agents offer a promising way to address these limitations (Chen et al., 2025).
+LLM agents have demonstrated the ability to role-play user personas, generate plausible sequences of decisions, and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024).
+Yet most prior work has focused on single-session tasks or constrained sandbox environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting their relevance for evaluating the dynamic and diverse behaviors seen on real-world websites.
+What remains missing is a scalable framework that can deploy large numbers of persona-driven agents into live web environments, allowing designers to test interface variants and obtain early behavioral signals before committing real user traffic.
+We address this gap by proposing Agent A/B (Fig.
+1), an end-to-end system for simulating large-scale online experiments using LLM-based agents.
+The system is designed to be interoperable, able to plug and play with existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)).
+As shown in Fig.
+1, Agent A/B integrates five coordinated components, spanning agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated post-testing analysis.
+At the core of this pipeline lies the LLM agent sub-system (Fig.
+3), which iteratively parses dynamic webpages into structured representations, reasons about the next action given its persona and intention, and grounds that action into concrete DOM operations on live sites.
+This design enables thousands of agents to run distributed sessions at scale, producing rich traces and statistical comparisons without manual intervention.
+Our position is that LLM agents should not replace real user testing; instead, they complement limited piloting sessions by offering earlier, faster, and lower-risk insights for design optimization.
+We demonstrate the system through a case study on Amazon.com, where we simulated a between-subjects A/B test comparing the existing full filter panel design against a reduced filter panel novel design powered by a similarity-based ranking algorithm.
+Using 1,000 agents (500 per condition), we observed that the reduced filter list led to more purchases (414 vs.
+403), a modest but statistically reliable increase.
+Average spending showed a small upward trend, though not statistically significant, and exploratory subgroup analyses suggested stronger effects for male and older personas.
+Importantly, these agent-based outcomes aligned directionally with the results of a parallel large-scale human A/B experiment, lending confidence to the system’s ability to surface meaningful behavioral signals.
+In summary, this work contributes:
+  1. (1)
+    Agent A/B, a system that deploys LLM-based agents to conduct scalable, persona-driven A/B testing on live webpages.
+  2. (2)
+    Empirical evidence from a case study on Amazon.com showing alignment between agent simulations and large-scale human A/B results.
+  3. (3)
+    Future design direction, demonstrating how agent-based simulations can extend A/B testing toward earlier prototyping, inclusive piloting, and automated design exploration.
 2. Related Work
@@
 The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites.
-[…]
-A timeline diagram illustrates the standard A/B testing workflow and highlights three key challenges.
-The timeline includes seven sequential phases: Feature Ideation (new UI layout or functionalities), A/B Testing Experiment Design (testing configuration and implementation), Data Monitoring (collect user interactions and rationales), Feature Deployment (final design decisions), Post-Testing Analysis (behavior patterns, failure cases, outcomes), Traffic Allocation (real user traffic toward variants), and Team Alignment (decision maker approval or rejection).
-The A/B Testing Period is marked across the Data Monitoring and Traffic Allocation stages.
-Three challenges are highlighted in red: (1) Scalable recruitment difficulties, (2) Limited testing opportunities, and (3) Slow user feedback loop.
-Each stage is connected along a central purple timeline.
+:
 2.2. Automated Experimentation and Interface Evaluation Tools
@@
 (2020) introduced CrowdUI, which used inverse reinforcement learning to infer task strategies from real user traces.
-With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human-like behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024).
+With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024).
 Park et al.
@@
 They called for innovations that support early-stage prototyping, internal evaluation without real user exposure, and predictive assessment of test outcomes before the actual A/B testing period starts.
-4. AgentA/B: Automated A/B Testing on the Web with LLM Agents
+4. Agent A/B
 Based on feedback from the formative study, we designed and developed Agent A/B a modular system that enables automated A/B testing on live websites with LLM agents to simulate realistic and diverse user interactions.
@@
 The system supports stratified analysis by agent demographics or personas to identify subgroup differences.
-It automatically computes effect sizes (e.g., absolute/relative changes in conversion-like behaviors) to aid interpretation.
 For instance, when testing redesigned filters, the system can reveal whether agents refined searches more, completed tasks faster, and purchased more, which offers early insights on usability and adoption risks before live deployment.
 Finally, Agent A/B maintains compatibility with external analytics pipelines.
-The result logs (in .JSON and .XSL) are exported in a format compatible with common data science tools, enabling downstream statistical modeling, significance testing, or integration with traditional A/B testing dashboards.
+The result logs (in .JSON and .XLS) are exported in a format compatible with common data science tools, enabling downstream statistical modeling, significance testing, or integration with traditional A/B testing dashboards.
 Users can use Agent A/B to analyze thousands of simulated interaction session results into actionable insights for interface evaluation.
@@
 (3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration.
-[…]
-A system diagram shows how a language model-based agent interacts with a web environment.
-The process begins with a web interface (Control/Treatment) feeding into an Environment Parsing Module, which extracts the webpage as a JSON structure and defines an action space (e.g., search, filter, purchase, product_click, terminate).
-This information, along with a predefined persona and intention, is sent to the LLM Agent (Action Prediction), which outputs the next action (e.g., search “solar filter for telescope”).
-This action is executed through the Action Execution Module, which sends it back to the Web Environment, completing the interaction loop.
+:
 4.2. Agent-Environment Interaction Architecture
@@
 The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B.
-In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).
-[…]
-A side-by-side comparison of two Amazon search results pages for the query ”solar filter for telescope.” Both pages show similar products but have differences in their filter options on the left sidebar.
-The left image displays additional filter categories under ”Department” such as ”Telescope Refractors,” ”Safety Goggles & Glasses,” and ”Telescope Eyepieces,” which are not present in the right image.
-Both images include nested subcategories under ”Camera & Photo Filters” and identical product listings on the right side of the page.
-This comparison highlights subtle differences in the user interface that could affect user navigation or filtering behavior.
+In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).:
 Action Execution Module.
@@
 If an action fails due to a missing selector or DOM mismatch, the system attempts fallback options such as retrying, scrolling into view, or re-parsing the page.
-Each successful or failed execution updates the environment and starts the next iteration of the loop.
+Each execution updates the environment and starts the next iteration of the loop.
 This interaction loop continues until the agent reaches a terminal condition.
@@
 5.1. Method: A/B Testing Scenario
-We designed a simulated A/B testing scenario focused on the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce.
-The goal was to evaluate how variations in filter configurations influence user behavior.
-{booktabs}
-colspec=lccc Control Condition
+We designed a simulated A/B testing scenario focused on a novel feature design of the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce.
+The goal was to evaluate how variations in filter configurations influence user behaviors such as dollar spending and product purchases.
+Control Condition
 Full Filter List
@@
 Average actions 15.96 6.05 6.60
-# of purchase * - 404 414
+# of purchase * - 403 414
 Average \$ spend - \$55.14 \$60.99
@@
 In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the users’ search query.
-Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions.
-Each agent was initialized with a persona profile and a shopping goal (e.g., “find a budget smart speaker under \$40 with strong customer reviews”).
-The allowed action space includes typical e-commerce flows: search, filter use, product examination, and cart actions.
+Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions, 500 assigned into the treatment condition and the other 500 assigned into the control condition.
+In parallel, we also conducted the real human online A/B testing with 1 million users randomly assigned into the control condition and another 1 million users in the treatment condition.
+Each agent was initialized with a persona profile and a shopping task (e.g., “find a budget smart speaker under \$40 with strong customer reviews”).
+The allowed action space for the LLM agent includes typical e-commerce flows: search, filter use, product examination, and cart actions.
 Persona generation followed the methodology in Chen et al.
@@
 Sessions ended either with task success (completion or no further predicted actions) or failure (e.g., looping behavior).
-We logged full action traces, metadata (e.g., duration, outcome), and agent rationales where applicable.
+We logged full action traces, metadata (e.g., duration, outcome), and the agent’s rationales where applicable.
 5.2. Finding: Alignment with Human Behavior
 To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using Agent A/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions.
-Table 5.1 presents the aggregated actions and statistics from sessions with LLM agents and with human participants.
-Human participants engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs.
+Table 2 presents the aggregated actions and statistics from sessions with the 1,000 LLM agents participating and with 2M human participants [^4] Due to company policy, the human participant’s behavior data in the treatment condition can not be shared..
+As shown in Table 2, human participants in the control condition engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs.
 6.05).
-They also clicked more products (6.96 vs.
+They also clicked on more product detail pages (6.96 vs.
 1.87) and performed more searches (6.40 vs.
@@
 This alignment supports the utility and validity of employing agent simulations for controlled UX evaluations, especially in scenarios where rapid interface assessment without live deployment is beneficial.
+Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.:
+(a) Average customer spending by gender groups across conditions.
+(b) Average customer spending by age groups across conditions.
+Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).:
 5.3. Findings: System Effectiveness Across Interface Variants
 To assess whether Agent A/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations.
-As shown in Table 5.1, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
+As shown in Table 2, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
 Agents in the treatment condition clicked on more products (2.09 vs.
@@
 0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability.
-Lastly, when looking at the final outcome of the sessions in the treatment condition and control condition, LLM agent simulated virtual customers purchased more products (414 v.s.
-404, $\chi^{2}(1)=5.51$, p=0.03) , and spent more money (\$60.99 v.s.
-\$55.14, t(998)=0.39, p=0.69) than the virtual customers in the control condition.
-Most of these comparisons, except the “# of product purchase” are not statistically significant, which suggests to the A/B test owner that this filter interface design demands a higher number of participants size (N).
-Importantly, the LLM-agent A/B testing results are consistent with the direction of changes in the human A/B testing results (which can not reported due to privacy policy).
-In summary, the A/B simulation conducted with AgentA/B reveals that LLM-based agents, when equipped with structured personas and embedded intentions, produce behaviorally rich and interface-sensitive interactions.
-These behaviors vary in predictable ways based on persona, align with human user baselines, and are capable of distinguishing between subtle interface design variants.
-These findings validate the system’s use as a behaviorally grounded, scalable alternative to conventional A/B testing, suitable for early-stage prototyping, pre-deployment design validation, and hypothesis-driven UX evaluation.
+When looking at the final outcomes of the sessions in the treatment and control conditions, LLM-agent simulated virtual customers purchased more products in the treatment condition compared to the control condition(414 v.s.
+404, $\chi^{2}(1)=5.51$, p=0.03).
+This represents a modest but statistically significant increase in purchase completion, suggesting that reducing the number of filter options improved the likelihood of customers making a purchase.
+Expressed in practical terms, the treatment condition resulted in roughly a 2–3% higher conversion rate, which could translate into millions of additional purchases per day on Amazon.
+In terms of average spending across all customers, LLM agent simulated virtual customers in the treatment condition spent somewhat more than those in the control condition (\$60.99 v.s.
+\$55.14).
+However, this difference was not statistically significant (t(998) = 0.39, p = 0.69).
+The mean difference of about \$6 is small relative to the variability in spending, suggesting that while the treatment may have had a slight positive effect on purchase amounts, substantially larger sample sizes would be required to reliably detect such an effect.
+Breaking down average spending by gender, the data show different patterns across groups.
+Female customers spent an average of \$49.18 in the control condition and \$51.30 in the treatment condition, indicating only a small increase under the treatment.
+Male customers, by contrast, showed a larger increase, from \$61.82 (control) to \$70.83 (treatment).
+While statistical tests were not conducted for each subgroup separately, the direction of change suggests that the simplified filter interface may have a stronger positive effect on male customers’ spending than on female customers.
+A clearer set of differences emerges when comparing average spending across age groups.
+Among younger customers ($\leq 35$ ), average spending actually decreased in the treatment condition compared to the control condition ($\mathdollar 56.68\rightarrow\mathdollar 43.41$).
+Middle-aged customers (35–55) spent more under treatment ($\mathdollar 59.21\rightarrow\mathdollar 71.19$), while older customers (¿55) showed the largest increase ($\mathdollar 40.24\rightarrow\mathdollar 79.07$).
+These results suggest that the reduced filter interface may be particularly beneficial for older customers, potentially because it helps reduce choice overload.
+Conversely, younger customers may rely more on detailed filtering to find products that match their preferences, leading to lower spending when filter options are reduced.
+In sum, reducing the number of filter options was associated with a higher purchase completion rate and possible increases in customer spending, though the latter effects did not reach statistical significance.
+Subgroup analyses suggest heterogeneity in response: males and older customers appear to benefit more from the simplified interface, while younger customers may prefer more filtering flexibility.
+Given the variability in spending outcomes, further testing with larger samples will be important to confirm these trends.
+Notably, the LLM-agent A/B testing results aligned directionally with patterns observed in human A/B testing .
+This convergence increases confidence that agent-based simulations can anticipate real user behaviors and provide early insight into interface performance.
+Overall, the A/B simulation conducted with Agent A/B shows that LLM-based agents, when guided by personas and embedded intentions, generate behaviorally rich and interface-sensitive interactions.
+These behaviors vary in predictable ways across personas, align with humans, and detect subtle differences in design.
+Such findings support the system’s role as a behaviorally grounded, scalable complement to conventional A/B testing.
+The results suggest Agent A/B is well-suited for early-stage prototyping, pre-deployment validation, and hypothesis-driven UX research.
 6. Discussion
@@
 In this section, we discuss the broader implications of our findings, system-level affordance, current limitations, and future research directions.
-6.1. Accelerating Design Iteration Through Simulated Feedback Loops
+6.1. Accelerating Design Iteration with Simulated Feedback
 One of the most immediate and impactful contributions of Agent A/B lies in its ability to expedite the feedback loop for A/B testing and design evaluation in web-based applications.
@@
 Appendix A Example of Persona Generated by Agent A/B
-{minted}
-[breaklines, breaksymbol=,breakanywhere]text Persona: Marcus
-Background: Marcus is a 35-year-old freelance graphic designer living in Austin, Texas.
+Persona: Marcus
+[…]
+Marcus is a 35-year-old freelance graphic designer living in Austin, Texas.
 After working for a decade in various creative agencies, he transitioned to freelancing to gain more control over his schedule and focus on passion projects, such as illustrating indie game assets and creating digital art for local musicians.
-Demographics:
-Age: 35
-Gender: Male
-Education: Bachelor’s degree in Visual Communication
-Profession: Freelance Graphic Designer
-Income: $70,000(variablebasedonprojects)\par FinancialSituation:Marcusearnsadecentlivingfromhisfreelancegigs,thoughhisincomecanfluctuate.He^{\prime}sfinanciallystablebutcautiousaboutbigexpenses.Hesetsasidepartofhisearningsfortravelandsoftwareupgrades,whichareessentialforhiswork.\par ShoppingHabits:Marcusenjoysdiscoveringuniqueornicheproducts,especiallytechgadgets,artsupplies,andstreetwear.Heprefersshoppingonlineforthevarietyandreadsreviewscarefully.He^{\prime}sbrand-loyalwhenitcomestotoolsherelieson,likehisdrawingtabletanddesignsoftware.Forclothing,he^{\prime}sdrawntobold,graphic-heavyitemsthatreflecthisartisticvibe.\par ProfessionalLife:Marcusworksfromahomestudiothatdoublesasacreativespace.Hecollaboratesremotelywithclientsfromdifferentindustries,jugglesmultipledeadlines,andoftenpullslatenights.Hefrequentlyupdateshisportfolioandmaintainsastrongsocialmediapresencetoattractnewclients.\par PersonalStyle:Marcushasanedgyandexpressivefashionsense.Hewearslarge-sizedclothingandgravitatestowarddarktoneswithpopsofneonorgraphicprints.Comfortisimportant,buthelikestomakeavisualstatementwithwhathewears.Hisgo-tooutfitisasofthoodiewithacustomdesign,blackjoggers,andhigh-topsneakers.\par\@add@PDF@RDFa@triples\par$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Marcus earns a decent living from his freelance gigs, though his income can fluctuate.
+Hes financially stable but cautious about big expenses.
+He sets aside part of his earnings for travel and software upgrades, which are essential for his work.
+[…]
+Marcus enjoys discovering unique or niche products, especially tech gadgets, art supplies, and streetwear.
+He prefers shopping online for the variety and reads reviews carefully.
+Hes brand-loyal when it comes to tools he relies on, like his drawing tablet and design software.
+For clothing, hes drawn to bold, graphic-heavy items that reflect his artistic vibe.
+[…]
+Marcus works from a home studio that doubles as a creative space.
+He collaborates remotely with clients from different industries, juggles multiple deadlines, and often pulls late nights.
+He frequently updates his portfolio and maintains a strong social media presence to attract new clients.
+[…]
+Marcus has an edgy and expressive fashion sense.
+He wears large-sized clothing and gravitates toward dark tones with pops of neon or graphic prints.
+Comfort is important, but he likes to make a visual statement with what he wears.
+His go-to outfit is a soft hoodie with a custom design, black joggers, and high-top sneakers.

### v4

@@
-Agent A/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents
+Agent A/B: Automated and Scalable A/B Testing on Live Websites with Interactive LLM Agents
 Abstract.
-A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long run times, and high operational cost.
-To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable and behavioral evidence before launch.
-In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases (414 vs.
-403), reproducing directional outcomes also observed in a parallel large-scale human experiment.
+A/B testing is central to UI/UX design, yet our formative study with six industry practitioners revealed that it is slowed by scarce user traffic, long runtimes, and high operational costs.
+To address these challenges, we introduce Agent A/B, an end-to-end system that deploys large language model (LLM) agents with structured personas to interact with live webpages and generate scalable behavioral evidence before launch.
+In a case study on Amazon.com, Agent A/B simulated a between-subjects A/B test of filter panel designs with 1,000 agents and found that the reduced filter list produced more purchases, reproducing directional outcomes also observed in a parallel large-scale human experiment.
 Results further suggest that agent-based simulations can detect interface-sensitive behavioral differences and surface subgroup patterns while offering faster, lower-risk insights.
@@
 large language models, LLM agents, user experience, user interface, A/B testing
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computingccs: Human-centered computing HCI design and evaluation methods
+journalyear: 2026copyright: ccconference: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems (CHI EA ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772363.3799039isbn: 979-8-4007-2281-3/2026/04ccs: Human-centered computing Usability testing
 Figure 1.
-Architecture of Agent A/B system.
-The system takes user input of agent and testing specifications to automatically 1) generate a large-scale of LLM agents, 2) prepare the testing conditions (e.g., traffic allocation), 3) perform agent-web interaction with the assigned web environment, 4) monitor and record agent behaviors as well as web interactions, and 5) conduct post-testing analysis so that the results will be returned to the end users.:
+Architecture of Agent A/B.
+From agent and test specifications, the system (1) generates LLM agents, (2, 3) runs controlled web interactions, and (4, 5) analyzes logged behaviors for experiment owners.
+A flowchart diagram illustrates the A/B testing workflow enabled by UXAgent.
+On the left, a system user (e.g., UX researcher or product manager) provides two types of input: LLM Agent specifications (e.g., number of agents, demographic distributions, attributes) and A/B Testing configuration (e.g., feature designs, agent intentions, metrics).
+These inputs lead to two parallel processes: LLM Agent Generation (Step 1) and Testing Condition Preparation (Step 2).
+The generated agents are then assigned to either a control or treatment group (Step 3), where they interact with their designated web environments.
+The control group uses the baseline interface, while the treatment group uses a modified (new) design.
+Simulation results are collected (Step 4) and passed to the Post-Testing Analysis module (Step 5), which returns A/B Testing Results (e.g., session outcomes, average actions, failure rates) back to the user.
+The visual includes labeled arrows, icons representing agents and interfaces, and text boxes describing each component.
 1. Introduction
-A/B testing—also known as online controlled experimentation—has become a cornerstone methodology for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015).
-By comparing alternative interface designs or feature functionalities with real users randomly assigned to control and treatment groups, A/B testing provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021).
-Major technology companies,including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), rely on hundreds of such tests each year to refine features, improve engagement, and guide strategic design choices.
-Despite its ubiquity, however, traditional A/B testing is often slowed by practical constraints, including heavy reliance on live user traffic, long turnaround times for statistically reliable results, and significant engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
-To better understand these challenges, we conducted a formative study with six industry practitioners who regularly run A/B tests in their day-to-day workflows.
-Their accounts highlight three recurring bottlenecks (Fig 2).
-First, there is a lack of lightweight methods for piloting design hypotheses before committing to large-scale traffic allocation.
-Second, user traffic itself is a scarce and contested resource, particularly for niche products or teams with competing priorities.
-Third, feedback cycles are slow: weeks may pass between experiment setup and actionable insights, discouraging exploration of alternative designs.
-Collectively, these constraints mean that many promising ideas are never tested, and early-stage prototyping rarely benefits from rigorous, behavioral-data grounded evaluation.
-Recent advances in large language model (LLM)–based agents offer a promising way to address these limitations (Chen et al., 2025).
-LLM agents have demonstrated the ability to role-play user personas, generate plausible sequences of decisions, and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024).
-Yet most prior work has focused on single-session tasks or constrained sandbox environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting their relevance for evaluating the dynamic and diverse behaviors seen on real-world websites.
-What remains missing is a scalable framework that can deploy large numbers of persona-driven agents into live web environments, allowing designers to test interface variants and obtain early behavioral signals before committing real user traffic.
-We address this gap by proposing Agent A/B (Fig.
-1), an end-to-end system for simulating large-scale online experiments using LLM-based agents.
-The system is designed to be interoperable, able to plug and play with existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)).
+A/B testing (i.e., online controlled experimentation) is a cornerstone for evaluating design decisions in modern web applications (Kohavi et al., 2009, 2007; Kohavi and Longbotham, 2015; Goswami et al., 2015).
+By comparing alternative interfaces or features with users randomly assigned to control and treatment groups, it provides rigorous, data-driven evidence for user experience optimization (Johari et al., 2017; Wang et al., 2019; Duan et al., 2021).
+Major technology companies, including Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017), run hundreds of such tests annually to refine features, improve engagement, and guide strategic design choices.
+Despite its ubiquity, traditional A/B testing is costly, slow to iterate, and difficult to revise when designs are flawed, due to long turnaround times and substantial engineering and organizational overhead (Kohavi et al., 2013; Xu et al., 2015; Fabijan et al., 2018).
+Our goal is to better support A/B testing experimenters in designing and running experiments.
+To ground our system design, we conducted a formative study with six industry practitioners who regularly run large-scale A/B tests.
+Their accounts reveal three recurring bottlenecks (Fig.
+2): (1) limited support for lightweight piloting before large-scale traffic allocation, (2) scarce and contested user traffic, and (3) slow feedback cycles between experiment setup and actionable insights.
+These constraints leave many promising ideas untested, and early-stage prototyping rarely benefits from rigorous, behaviorally grounded evaluation.
+Recent advances in large language model (LLM) agents offer a promising way to address these limitations (Chen et al., 2025b; Zhang et al., 2024; Jia et al., 2024; Zhang et al., 2025b; Chen et al., 2025a; Wang et al., 2025a, b; Zhang et al., 2025a; Wang et al., 2026).
+LLM agents can role-play user personas and perform multi-step interactions in domains such as e-commerce (Yao et al., 2022), booking (Zhou et al., 2023), and information search (Koh et al., 2024).
+However, most prior work focuses on single-session tasks or constrained environments (Li et al., 2024; Luo et al., 2023; Park et al., 2023; Hua et al., 2023; Wu et al., 2023; Yao et al., 2023), limiting applicability to real-world websites.
+What remains missing is a scalable framework that deploys large numbers of persona-driven agents in live web environments to surface early behavioral signals before allocating real user traffic.
+We address this gap with Agent A/B (Fig.
+1), an end-to-end system for simulating large-scale online experiments with LLM-based agents.
+The system is interoperable and can plug into existing agent stacks (e.g., Claude computer-use agents or ReAct (Yao et al., 2023)).
 As shown in Fig.
-1, Agent A/B integrates five coordinated components, spanning agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated post-testing analysis.
-At the core of this pipeline lies the LLM agent sub-system (Fig.
-3), which iteratively parses dynamic webpages into structured representations, reasons about the next action given its persona and intention, and grounds that action into concrete DOM operations on live sites.
-This design enables thousands of agents to run distributed sessions at scale, producing rich traces and statistical comparisons without manual intervention.
-Our position is that LLM agents should not replace real user testing; instead, they complement limited piloting sessions by offering earlier, faster, and lower-risk insights for design optimization.
-We demonstrate the system through a case study on Amazon.com, where we simulated a between-subjects A/B test comparing the existing full filter panel design against a reduced filter panel novel design powered by a similarity-based ranking algorithm.
-Using 1,000 agents (500 per condition), we observed that the reduced filter list led to more purchases (414 vs.
-403), a modest but statistically reliable increase.
-Average spending showed a small upward trend, though not statistically significant, and exploratory subgroup analyses suggested stronger effects for male and older personas.
-Importantly, these agent-based outcomes aligned directionally with the results of a parallel large-scale human A/B experiment, lending confidence to the system’s ability to surface meaningful behavioral signals.
+1, Agent A/B integrates five components: agent specification, testing configuration, live agent–web interaction, behavioral monitoring, and automated analysis.
+At its core, LLM agents iteratively interpret dynamic webpages and ground persona-driven decisions into concrete DOM actions, enabling thousands of distributed sessions and scalable behavioral comparisons without manual intervention.
+We emphasize that LLM-agent-based A/B testing is not a replacement for real user testing, but a complementary tool to help experiment owners (e.g., UX researchers and product managers) mitigate traffic scarcity, slow iteration cycles, and collaboration challenges (Kuang et al., 2022; Feng et al., 2023).
+We demonstrate Agent A/B through a case study on Amazon.com, simulating a between-subjects A/B test comparing the existing full filter panel with a reduced, similarity-based design.
+Using 1,000 agents (500 per condition), the reduced design produced more purchases than the baseline, with a modest but statistically significant increase.
+Spending showed a small upward trend, and agent outcomes aligned directionally with a parallel large-scale human A/B experiment, suggesting the system can surface meaningful behavioral signals.
 In summary, this work contributes:
   1. (1)
-    Agent A/B, a system that deploys LLM-based agents to conduct scalable, persona-driven A/B testing on live webpages.
+    Agent A/B, an end-to-end system for scalable, persona-driven A/B testing with LLM-based web agents.
   2. (2)
-    Empirical evidence from a case study on Amazon.com showing alignment between agent simulations and large-scale human A/B results.
+    Empirical evidence from an Amazon.com case study showing alignment between agent simulations and human A/B results.
   3. (3)
-    Future design direction, demonstrating how agent-based simulations can extend A/B testing toward earlier prototyping, inclusive piloting, and automated design exploration.
+    Design implications for using agent-based simulation to support early prototyping and exploratory experimentation.
 2. Related Work
-We situate our work at the intersection of A/B testing, user behavior modeling, autonomous agents, and HCI research on interactive prototyping and simulation.
-We review four relevant threads of work: (1) limitations of traditional A/B testing practices, (2) advances in automated and optimized experimentation workflows, (3) LLM-based agent simulation across domains, and (4) interactive agents in web environments.
-2.1. Limitations of Traditional A/B Testing in Practice
-A/B testing has become a foundational methodology of comparing two versions of interfaces and functionalities of a webpage, app, or other digital assets to enable data-driven decisions about design strategies based on user behavior differences (Kohavi et al., 2009, 2007).
-The application of A/B testing spans both industry and academia, where organizations (e.g., major companies like Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017)) routinely rely on A/B testing to streamline the design, development, and deployment of products.
-However, numerous studies in HCI communities have identified structural limitations in this widely adopted methodology.
-For instance, Fabijan et al.
-(2017) conducted an in-depth study across software-intensive organizations and found that A/B testing introduces slow iteration loops, costly feature development, and high failure rates—particularly when hypothesis formulation is weak.
-These findings were reinforced in a broader empirical survey by Fabijan et al.
-(2018), who emphasized the lack of actionable early-stage insights during the experimentation process.
-These works collectively underscore the HCI community’s recognition that while A/B testing provides rigor and promising benefits, it falls short in flexibility, speed, and insight generation.
-Figure 2.
-The workflow of web A/B testing and three challenges reported from formative study: 1) the cost and difficulty of securing considerable user traffic for significant results, 2) the whole A/B testing period can span across weeks and months, and 3) limited testing opportunites.
-:
-2.2. Automated Experimentation and Interface Evaluation Tools
-Given the aforementioned limitations, the HCI community has grown significant interest in systems that accelerate interface experimentation (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022; Swearngin et al., 2012).
-For example, Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) allowed designers to rapidly prototype for physical interaction such as touch interfaces.
-Fuse (Kuznetsov et al., 2022) enabled the rapid creation of context-aware UI behaviors from demonstrations.
-On the experimentation side, Gilotte et al.
-(2018) proposed techniques for offline A/B testing, using logged user interaction data to estimate counterfactual outcomes for unobserved experimental variants.
-This reduces the need for live deployments but depends heavily on rich user logs and pre-existing infrastructure.
-Tamburrelli and Margara (2014) reframed A/B testing as a search problem within a design space, suggesting that evolutionary algorithms could be used to automate variant generation.
-These systems aim to reduce human effort in running and prioritizing experiments, yet still rely on historical data or user deployment.
-In contrast, our work explores agent-driven simulation as a complementary mechanism to automatically evaluate designs without live user data.
-2.3. User Behavior Simulation: From Cognitive Models to LLM Agents
-A longstanding tradition in HCI focuses on simulating user behavior through cognitive models such as GOMS and ACT-R (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996; Gray and Altmann, 2001).
-Building upon these traditional approaches, a number of works have leveraged inverse reinforcement learning (IRL) to infer user intentions from observed behaviors.
-Ziebart et al.
-(2009) introduce a probabilistic approach to IRL, enabling the modeling of complex user strategies in dynamic environments.
-However, these models are labor-intensive, require domain expertise, and do not generalize easily to open-ended tasks like online shopping.
-More recent work uses data-driven models to simulate user behavior in online environments.
-For example, Paranjape et al.
-(2016) used server logs to reconstruct navigation paths and optimize hyperlink structures.
-Oppenlaender et al.
-(2020) introduced CrowdUI, which used inverse reinforcement learning to infer task strategies from real user traces.
-With the rise of large language models (LLMs), researchers have begun to explore how LLMs can simulate human behaviors across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024).
-Park et al.
-(2023) built a generative society of agents simulating daily social behavior in a simulated town.
-Horton (2023) evaluated whether LLMs could replicate outcomes from real behavioral experiments on prosociality.
-Furthermore, Lu et al.
-(2025a) demonstrates that fine-tuning large language models (LLMs) on real-world behavioral data significantly enhances their ability to generate accurate user actions.
-These systems suggest that LLMs can produce high-fidelity behavioral simulations across diverse tasks.
-Our work extends this direction into the visually rich, dynamic domain of live web interfaces, focusing on shopping behavior and automated A/B testing experiments.
-2.4. LLM Agents in Web Environments
-A growing body of work has investigated how autonomous agents can operate within web-based environments, driven by advancements in both machine learning and HCI communities.
-For instance, WebShop (Yao et al., 2022) introduced a shopping benchmark with templated webpages for studying goal-directed navigation.
-WebArena (Zhou et al., 2023) extended this to multi-domain service tasks and became a standard evaluation for web agent development.
-VisualWebArena (Koh et al., 2024) incorporated vision-language models to parse visual cues in interface design.
-Similarly, WebVoyager (He et al., 2024) and WILBUR (Lutz et al., 2024) emphasized task generalization and adaptive memory in agents interacting with open-ended web content.
-However, all these systems operate within simulated environments, which, while useful for reproducibility, abstract away many complexities of the real webpage, such as dynamic loading, layout noise, unexpected modals, and inconsistent structure.
-Furthermore, the focus of existing work is primarily on task completion, not on evaluating or comparing the user experience across different design configurations.
-This gap reveals a growing need for systems that can adapt to the diversity and complexity of real-world web environments in which designers often have to balance trade-offs between design constraints and usability.
-Addressing this gap requires new system architectures and evaluation frameworks that can support interaction in the wild and enable comparative analysis of design decisions in terms of agent efficiency, robustness, and user-centered outcomes.
-Participant Gender Job Title Team
-P1 F Project Manager PM Team
-P2 M Software Development Manager Engineer Team
-P3 F Product Manager PM Team
-P4 M Product Manager PM Team
-P5 M Product Manager PM Team
-P6 M Machine Learning Researcher Science Team
-Table 1. Demographics of formative study participants.
+We situate our work at the intersection of A/B testing, user behavior modeling, and HCI research on interactive prototyping and simulation.
+We review two threads: (1) limitations of traditional A/B testing and advances in automated experimentation workflows, and (2) LLM-based agent simulation across domains.
+2.1. Limitations and Automation of A/B Testing
+A/B testing is a foundational methodology for comparing interface and feature variants to support data-driven design decisions (Kohavi et al., 2009, 2007), and is widely used in both industry and academia, including by companies such as Microsoft (Machmouchi et al., 2017), Yahoo (Tagami et al., 2014), Netflix (Amatriain, 2013), and Amazon (Hill et al., 2017).
+Despite its ubiquity, prior HCI research documents persistent limitations, including slow iteration cycles, high development costs, limited early-stage feedback, and methodological challenges such as metric selection, data quality, and statistical reliability under real-world constraints (Fabijan et al., 2017, 2018; Gupta et al., 2019; Larsen et al., 2022).
+Beyond methodological issues, experimentation is shaped by socio-technical factors: UX practitioners face collaboration overhead, fragmented analyses, time pressure, and limited influence over data and model decisions in AI-intensive products (Kuang et al., 2022; Feng et al., 2023; Zdanowska and Taylor, 2022; Lu et al., 2022).
+In response, the HCI community has explored tools for rapid prototyping, in-situ sensemaking, and offline or automated experimentation (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022; Oppenlaender et al., 2020; Gilotte et al., 2018; Tamburrelli and Margara, 2014), but these approaches still rely on historical data or live user traffic.
+Our work builds on these efforts by investigating agent-driven simulation as a complementary approach for evaluating interface designs without consuming real user traffic.
+2.2. User Behavior Simulation with LLM Agents
+HCI has long studied user behavior simulation through cognitive models such as GOMS and ACT-R (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996; Gray and Altmann, 2001), and later through inverse reinforcement learning to infer user intentions from observed behavior (Ziebart et al., 2009; Kangasrääsiö and Kaski, 2018).
+While these approaches enable fine-grained modeling, they are labor-intensive, require domain expertise, and do not generalize well to open-ended tasks like online shopping; data-driven methods using server logs improve scalability but still rely on historical traces and fixed environments (Paranjape et al., 2016).
+Recent advances in large language models (LLMs) have renewed interest in behavior simulation across complex domains (Ge et al., 2023; Wan et al., 2024; Wu et al., 2024), with prior work demonstrating LLM agents’ ability to simulate social behavior (Park et al., 2023), reproduce behavioral experiments (Horton, 2023), Agent-as-a-judge evaluation (Chen et al., 2025c), and improve action fidelity when trained on real-world data (Lu et al., 2025).
+In parallel, web-agent benchmarks such as WebShop (Yao et al., 2022), WebArena (Zhou et al., 2023), VisualWebArena (Koh et al., 2024), WebVoyager (He et al., 2024), and WILBUR (Lutz et al., 2024) have advanced autonomous web interaction, but remain largely confined to simulated or templated environments and focus on task completion rather than comparative user experience evaluation.
+Our work bridges these lines by deploying persona-driven LLM agents directly on live web interfaces to enable scalable user behavior simulation for A/B testing and early-stage design evaluation.
 3. Formative Study: Understanding Challenges in A/B Testing Workflows
 3.1. Formative Study Method
-To inform the design of the LLM agent-based A/B testing system, we conducted a formative study with six professionals experienced in designing and running A/B tests in industry settings.
-Through semi-structured interviews, we asked participants to recount a recent A/B test: how it began and ended, the stages involved, technologies used, and challenges encountered.
-We also probed how they addressed issues and why those strategies were chosen.
-These interviews helped us understand current practices, challenges, and user needs, which later informed our design decisions for the system.
-We used a snowball sampling strategy (Biernacki and Waldorf, 1981) to recruit participants who currently work in the e-commerce industry and have experience designing and conducting A/B tests.
-We started with our friends, colleagues, and connections who fit the selection criteria and then asked them to refer their connection to participate in our study.
-All interviewees were located in the United States; A total of 6 participants (four product managers, one software development manager, and one machine learning researcher) were recruited to participate in the interview study (see Table 1 for details).
-Interviews were conducted remotely, with audio recorded upon consent.
-Sessions lasted 45–62 minutes.
-We transcribed and de-identified all recordings before conducting a grounded theory analysis (Muller and Kogan, 2012).
-Two co-authors independently performed open coding on practices, tools, and challenges.
-Through collaborative discussion, we developed a shared codebook and organized codes into categories (e.g., experiment design, system bottlenecks), which were then synthesized into broader themes such as the A/B testing lifecycle and opportunities for system support.
-We finally applied the developed set of codes to the whole corpus of transcripts.
+To inform the design of our LLM agent-based A/B testing system, we conducted a formative study with six industry professionals experienced in designing and running large-scale A/B tests.
+We used semi-structured interviews to elicit recent A/B testing experiences, focusing on workflows, tools, challenges, and how participants addressed them.
+Participants were recruited via snowball sampling (Biernacki and Waldorf, 1981) from the e-commerce industry until thematic saturation.
+All six participants were based in the United States (four product managers, one software development manager, and one machine learning researcher) and routinely managed A/B tests with over one million users.
+Interviews were conducted remotely, lasted 45–62 minutes, and were audio-recorded with consent.
+Recordings were transcribed, de-identified, and analyzed using grounded theory (Muller and Kogan, 2012).
+Two authors independently coded the data, reconciled differences, and synthesized themes related to A/B testing workflows, bottlenecks, and opportunities for system support.
 3.2. Formative Study Findings
-In this section, we first report the overall workflow of an A/B testing project lifecycle (Fig.2), followed by the challenges and coping strategies reported by industrial practitioners.
-We conclude this section with design requirements for an LLM-Agent based simulation system for A/B testing.
 A/B Testing Project Lifecycle
-The A/B testing project life cycle, as reported in the interview data, unfolds across seven interdependent stages (Fig.2).
-It begins with new feature ideation, where ideas originate from individuals or teams and can range from minor UI changes to novel functionalities.
-Once proposed, the idea enters a team alignment and buy-in stage, involving collaborative discussion and refinement, culminating in approval from higher-level decision-makers.
-With alignment secured, the A/B testing experiment design stage follows, where critical aspects such as user segmentation, control/treatment conditions, success metrics, and evaluation criteria are meticulously planned—this stage precedes any development to ensure methodological rigor.
-Next, the feature deployment and iteration stage involves cross-functional collaboration to build and refine the feature, often demanding the most time and resources.
-Upon completion, the A/B test experiment launch integrates the feature into a live A/B testing environment where user exposure is randomized and monitored.
-After the test concludes, post-experiment data analysis is conducted with data scientists to assess whether success criteria were met and to understand the underlying reasons.
-Finally, the feature decision is made based on the outcome: successful features are launched, while unsuccessful ones are shelved.
-Participants reported that the whole process can take anywhere from 3 months to a year.
-As aforementioned, the A/B test process is definitely a team sport, where different job roles need to collaborate with each other to design and develop the feature, and to design, launch, and analyze the A/B test experiment.
-Challenges and Coping Strategies
-Interviewees described several challenges in the current A/B testing workflow.
-One major issue is the high cost of new feature development, which often requires multiple engineers working for months to produce a functional version.
-During this lengthy process, it’s difficult to gather early user feedback to iterate on the design.
-The only formal feedback comes through the A/B test experiment, which may be too late.
-Some feature owners use colleagues as alpha testers, but this feedback is typically biased.
-Thus, interviewees expressed a need for lightweight prototyping and early-stage user testing methods that can offer feedback before committing to a formal A/B test.
-“We had features … required changes across all [UI] stacks and beyond in other services, and it took one year and a half [to develop]…” (P2, Software Development Manager)
-Another common challenge is competing for user traffic.
-When multiple teams want to test features that affect similar UI components, their experiments must be serialized, as running them in parallel would interfere with results.
-Currently, prioritization happens late in the process after development and just before the test launch, which could significantly delay the following steps.
-Interviewees called for a more data-driven way to manage and prioritize experiment launches.
-“The biggest pain point, I think, is ‘traffic lanes’ … So experiments look at the same area; you cannot run them in parallel, so we have to split the [user] traffic.
-But we don’t have an automated way to split the traffic properly… ” (P2, Software Development Manager)
-Some teams reported that up to one-third of experiments fail to meet their predefined success criteria.
-This high failure rate is frustrating because A/B testing in its current form leads to a one-shot decision: if a feature misses even one metric, the opportunity is lost.
-Interviewees expressed interest in predictive tools that could help forecast experiment outcomes so that they can revise features or experimental designs before the launch.
-“The hypothesis behind [a failed feature design] that is it’s easier to access [the UI element when put it at the bottom] … and to match our competitors.
-[A/B test experiment shows] customer hates it, [because] they can no longer find the UI element that they used to and they quit.
-We saw the abandonment rate go through the roof [in the A/B test experiment].” (P3, Product Manager)
-In summary, while A/B testing is widely adopted and supported by infrastructure such as user-routing platforms, interviewees identified several shortcomings: the high development cost, the lack of user traffic prioritization, and frequent experiment failure.
-They called for innovations that support early-stage prototyping, internal evaluation without real user exposure, and predictive assessment of test outcomes before the actual A/B testing period starts.
+Interview data revealed a seven-stage A/B testing lifecycle (Fig.
+2).
+The process begins with feature ideation, where teams propose new UI changes or functionalities, followed by team alignment and buy-in involving discussion, refinement, and managerial approval.
+Teams then conduct experiment design, specifying user segmentation, control and treatment conditions, and success metrics prior to development.
+The feature development and iteration stage follows, requiring cross-functional collaboration and substantial engineering effort.
+Once implemented, the experiment launch deploys the feature into a live A/B testing environment with randomized user exposure.
+Afterward, post-experiment analysis interprets outcomes and behavioral drivers, leading to a final feature decision to launch or discard the feature.
+Participants reported that this end-to-end process typically spans three months to over a year and requires close coordination across multiple roles.
+Challenges and Coping Strategies.
+Interviewees highlighted recurring challenges in current A/B testing workflows, including the high cost and long timelines of feature development, which leave little opportunity for early user feedback before formal testing.
+To cope, some teams rely on internal alpha testing with colleagues, though participants noted this feedback is often biased and insufficient for guiding iteration, underscoring the need for lightweight prototyping and early-stage evaluation.
+Interviewees also described intense competition for user traffic, which forces experiments affecting similar interface components to be serialized and prioritized late in the process, delaying deployment and iteration.
+Finally, high experiment failure rates often result in one-shot decisions with limited opportunities for refinement, leading participants to call for tools that enable internal evaluation, data-driven prioritization, and predictive assessment of experiment outcomes prior to live A/B testing.
 4. Agent A/B
-Based on feedback from the formative study, we designed and developed Agent A/B a modular system that enables automated A/B testing on live websites with LLM agents to simulate realistic and diverse user interactions.
-Traditional A/B testing is bottle-necked by a number of long-standing limitations, including the expensive requisites of large sample sizes and user traffic for significant results, the complicated and prolonged procedures for experiment preparation, execution, and analysis, as well as the limited resources that can only support experiments with very few designs/features.
-However, our system takes advantage of LLM agents for low-cost, rapid feedback loop, and high-throughput evaluation of interface variants without relying on live user deployment.
-The system is designed with modularity and extensibility in mind, which can be accommodated in different web environments, target user populations, and LLM agent architectures.
-In this section, we describe the Agent A/B system architecture in detail, including the end-to-end A/B testing pipeline, the interaction loop of LLM agent and the web environment, and the implementation techniques that support robust and high-fidelity simulation possible on live web platforms.
 4.1. System Overview and Pipeline
+Agent A/B is an end-to-end system for simulating A/B tests on websites with persona-driven LLM agents.
+Experiment owners provide (1) agent specifications (e.g., number of agents and persona attributes), (2) A/B testing configurations (e.g., task intentions and behavioral metrics), and (3) two functional web variants for the control and treatment conditions (Fig.
+1).
+Given these inputs, Agent A/B orchestrates four coordinated modules: (i) LLM Agent Generation, (ii) Testing Preparation, (iii) Autonomous A/B Simulation, and (iv) Post-Testing Analysis.
+First, the LLM Agent Generation Module instantiates a population of agents by generating diverse personas and pairing them with the user-specified task intentions, ensuring variability in demographics and behavioral tendencies while adhering to the experiment constraints.
+The experiment owner specifies a target demographic distribution and provides an example persona as a style reference; a persona pool is initialized with that example.
+The system then iteratively samples an existing persona from the pool, samples demographic attributes from the target distribution, and prompts the LLM to produce a new persona matching both the sampled demographics and the reference style.
+Each generated persona is added back to the persona pool, enabling diverse yet stylistically consistent personas at scale.
+Second, the Testing Preparation Module assigns agents to control and treatment conditions via traffic allocation and checks balance across key persona attributes, re-splitting when needed to reduce distributional skew.
+Third, in the Autonomous A/B Simulation Module, agents interact with their assigned web variant in isolated browser sessions using a recurrent perceive–decide–act loop (Section 4.2), while the system records step-level traces of agent actions and web states.
+Finally, the Post-Testing Analysis Module aggregates traces across conditions to compute outcome metrics (e.g., completion or purchase rates), interaction efficiency measures (e.g., actions per session, session duration), and behavior patterns (e.g., search or filter usage), and supports stratified analyses by persona attributes to surface subgroup differences.
+Together, these modules enable scalable comparative evaluation of interface variants and provide early behavioral signals without relying on live user traffic.
+Full details of the system design can be found in Appendix A.
+Control Condition
+Full Filter List
+Human, N=1M Control Condition
+Full Filter List
+Agent, N=500 Treatment Condition
+Reduced Filter List
+Agent, N=500
+Search 6.40 1.42 1.43
+Click_product 6.96 1.87 2.09
+Click_filter_option 0.33 0.58 0.60
+Purchase 0.62 0.81 0.83
+Stop - 0.19 0.17
+Average actions 15.96 6.05 6.60
+# of purchase * - 403 414
+Average \$ spend - \$55.14 \$60.99
+Table 1.
+Comparison of human customers’ actions per session in control condition (current design with full filter list), and the LLM agents as virtual customers in both control condition and treatment condition (new design with reduced filter list).
+The LLM agents in treatment condition purchased significantly more items than the ones in the control condition.
+$\chi^{2}(1)=5.51$, p-value ¡ 0.05
+4.2. Agent-Environment Interaction Architecture
+At the core of Agent A/B is an iterative interaction loop in which each LLM agent operates directly on a live web environment and continuously adapts its actions based on evolving page state.
+This loop consists of three tightly coupled components: an Environment Parsing Module, an LLM Agent, and an Action Execution Module (Fig.
+3).
+Together, they enable robust and scalable interaction with dynamic, real-world webpages.
+At each step, the Environment Parsing Module extracts the current webpage into a structured representation of task-relevant elements (e.g., products, filters, prices) and the available action space, reducing noise from raw HTML or visual clutter.
+This observation, together with the agent’s persona, intention, and interaction history, is passed to the LLM Agent, which performs language-based reasoning and planning to select the next structured action (e.g., search, click filter, select item, purchase), independent of any specific agent architecture.
+The Action Execution Module then translates this action into browser-level interactions and executes it in the live environment, incorporating basic fault handling such as retries or re-parsing to accommodate dynamic page changes.
+This perceive–decide–act loop continues until task completion or failure, while Agent A/B logs fine-grained traces of page states, actions, and outcomes for downstream A/B-style analysis.
+5. Case Study: A/B Testing with Reduced Filter Options on Amazon.com
+We evaluate Agent A/B through a simulated A/B testing case study on Amazon’s shopping interface, focusing on a redesign of the left-side filtering panel, a common target for UI experimentation in e-commerce.
+The treatment condition replaces the full filter list with a similarity-based ranking design that suppresses low-relevance options, while the control condition retains the existing interface.
+Using Agent A/B, we generated a large pool of agent personas and sampled 1,000 agents (500 per condition) to simulate shopping sessions under realistic demographic distributions and task intentions, in parallel with a large-scale human A/B test on the same feature.
+Agents interacted with the live website using standard e-commerce actions, and the system logged full interaction traces and outcomes for analysis.
+Full details of the scenario design, agent setup, infrastructure, and cost analysis are provided in Appendix B.
+5.1. Finding: Alignment with Human Behavior
+To assess how well Agent A/B aligns with real user behavior, we compared a simulated A/B test using 1,000 LLM agents with results from a parallel online A/B experiment involving 2M human participants under identical task conditions (Table 1).
+Humans exhibited more exploratory interaction patterns, including longer sessions and more frequent searches and product views, while agents followed more goal-directed trajectories with fewer actions.
+Despite these differences in interaction style, agents and humans showed comparable purchase rates and similar use of filter options, indicating alignment in decision outcomes for intention-driven tasks.
+Overall, these results suggest that while LLM agents simplify exploratory behavior, they capture key decision-making signals relevant for early-stage UX evaluation, supporting the use of agent-based simulation as a complementary tool to human A/B testing.
+5.2. Findings: System Effectiveness Across Interface Variants
+To evaluate whether Agent A/B can detect differences between interface variants, we compared LLM agent behavior across control and treatment conditions with different filter designs.
+As shown in Fig.
+6 and Table 1, agents in the treatment condition interacted more with products and filters and completed more purchases than those in the control condition (414 vs.
+404; $\chi^{2}(1)=5.51$, $p=0.03$), indicating a modest but statistically reliable increase.
+Average spending exhibited a small upward trend under the treatment design, though this effect did not reach statistical significance.
+Subgroup analyses further revealed heterogeneous responses across personas (Fig.
+6): agents representing older and male customers showed larger increases in spending under the simplified filter design, while younger agents exhibited decreased spending, suggesting differing preferences for filter complexity.
+Importantly, these agent-based patterns aligned directionally with results from human A/B testing, supporting Agent A/B ’s ability to surface meaningful behavioral signals and detect subtle interface effects prior to live deployment.
+6. Discussion
+Our design and evaluation of Agent A/B demonstrate how agent-based simulation can serve as a design support mechanism in the web interface design lifecycle.
+Rather than replacing real-user experimentation, Agent A/B complements existing workflows by enabling earlier, faster, and lower-risk exploration, aligning with broader trends in HCI that leverage large-scale simulation for iterative design and decision-making (Park et al., 2023; Horton, 2023; Gao et al., 2024).
+Simulation-Supported Design Iteration and Exploration.
+A key contribution of Agent A/B is its ability to accelerate design iteration by enabling agent-based pilot experiments prior to A/B testing.
+Although A/B testing remains the gold standard for design validation, it is constrained by limited user traffic, long feedback cycles, and substantial engineering and organizational overhead (Fabijan et al., 2017; Kohavi et al., 2013; Xu et al., 2015).
+By supporting rapid evaluation of interface variants through agent simulations, Agent A/B allows designers to obtain early signals without implementing and deploying their new design to real users.
+This capability extends prior prototyping and experimentation tools (Lasecki et al., 2015; Hartmann et al., 2006; Kuznetsov et al., 2022) by grounding early feedback in end-to-end interaction traces, and enables simulation-driven methodologies such as parallel evaluation and computational design exploration (Kumar et al., 2013).
+Inclusive Piloting and Risk-Aware Testing.
+Beyond accelerating iteration, Agent A/B supports inclusive and risk-aware piloting for user populations that are difficult, costly, or ethically sensitive to recruit.
+Groups such as older adults or users with lower digital literacy are often underrepresented in early experimentation, even though design flaws may disproportionately affect them.
+By configuring agent personas to reflect diverse demographic and behavioral characteristics, designers can explore how features may impact different user groups before exposing real users to potential harm.
+This capability complements prior crowd-based and proxy evaluation approaches (Bernstein et al., 2011; Oppenlaender et al., 2020) by offering an automated alternative that helps surface subgroup disparities early, enabling more robust and informed A/B testing, and aligns with emerging evaluation techniques using agent judges (Chen et al., 2025c).
+7. Conclusion
+In this paper, we presented Agent A/B, a system that enables large-scale, LLM agent-based simulation of A/B testing for web interfaces.
+Our evaluation demonstrated that LLM agents exhibit realistic, goal-aligned behaviors, and are sensitive to interface variations.
+By supporting rapid, risk-free behavioral piloting, Agent A/B introduces a new phase for agent-based piloting in the design life-cycle that complements traditional A/B testing and expands the scope of early-stage UX evaluation.
+We envision future extensions that further enhance agent fidelity, broaden domain coverage, and integrate simulation into intelligent design optimization workflows.
+Appendix A Agent A/B
+Figure 2.
+The workflow of web A/B testing and three challenges reported from formative study: (1) the cost and difficulty of securing considerable user traffic for significant results, (2) the whole A/B testing period can span across weeks and months, and (3) limited testing opportunites.
+A timeline diagram illustrates the standard A/B testing workflow and highlights three key challenges.
+The timeline includes seven sequential phases: Feature Ideation (new UI layout or functionalities), A/B Testing Experiment Design (testing configuration and implementation), Data Monitoring (collect user interactions and rationales), Feature Deployment (final design decisions), Post-Testing Analysis (behavior patterns, failure cases, outcomes), Traffic Allocation (real user traffic toward variants), and Team Alignment (decision maker approval or rejection).
+The A/B Testing Period is marked across the Data Monitoring and Traffic Allocation stages.
+Three challenges are highlighted in red: (1) Scalable recruitment difficulties, (2) Limited testing opportunities, and (3) Slow user feedback loop.
+Each stage is connected along a central purple timeline.
+Figure 3.
+One action prediction iteration of the automated web testing in Agent A/B with an LLM agent.
+(1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session.
+In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces.
+(3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration.
+The light-weight Environment Parsing Module and Action Execution Module are web-site specific by their nature, and other modules include LLM Agent are generalizable to different websites.
+A system diagram shows how a language model-based agent interacts with a web environment.
+The process begins with a web interface (Control/Treatment) feeding into an Environment Parsing Module, which extracts the webpage as a JSON structure and defines an action space (e.g., search, filter, purchase, product_click, terminate).
+This information, along with a predefined persona and intention, is sent to the LLM Agent (Action Prediction), which outputs the next action (e.g., search “solar filter for telescope”).
+This action is executed through the Action Execution Module, which sends it back to the Web Environment, completing the interaction loop.
+A.1. System Overview and Pipeline
 Agent A/B is designed as an end-to-end simulation system for LLM agent-based A/B testing in live browser environments.
-Analogous to the preparation stage of real user A/B testing, the end users of our system (i.e., UX researchers or product managers) need to determine details of A/B testing designs and provide two web environment variants to be tested.
+Analogous to the preparation stage of real user A/B testing, the experiment owners (i.e., UX researchers or product managers) need to determine details of A/B testing designs and provide two web environment variants to be tested.
 After taking the user inputs, the system operates on four LLM-powered modules (Fig.
@@
 In our e-commerce scenario, the tested design feature is the layout of the filter panel on the website.
-Once these inputs are provided by end users, the LLM Agent Generation Module queries the backend LLM to generate the specified number of LLM agents with diverse persona descriptions and intentions.
+Once these inputs are provided by the experiment owners, the LLM Agent Generation Module queries the backend LLM to generate the specified number of LLM agents with diverse persona descriptions and intentions.
 The query explicitly instructs the LLM to ensure the generated LLM agent persona and intentions must follow the user-provided agent specifications.
 In our experiment in Section 5, we leveraged this module to generate 100,000 agents.
-We provide a sample persona generated by our system in Appendix A.
-After generating the LLM agents as the candidate A/B testing participants, the Testing Preparation Module performs the agent traffic allocation by splitting the agents into control (without new feature) and treatment (with new feature) groups, and each group is assigned to interact with the corresponding web environments.
-The statistics of agent characteristic distributions will be calculated within each group to ensure that the distribution of the LLM agents is relatively balanced; if the statistics are not balanced, the Testing Preparation Module will re-execute agent traffic allocation until satisfactory.
-These web environments for both groups need to be launched using independent browser instances controlled through ChromeDriver (for web environment parsing) and Selenium WebDriver (for automated interaction execution) integration.
+We provide a sample persona generated by our system in Appendix C.
+After generating the LLM agents as the A/B testing participants, the Testing Preparation Module performs the agent traffic allocation by splitting the agents into control (without new feature) and treatment (with new feature) groups, and each group is assigned to interact with the corresponding web environments.
+The statistics of agent characteristic distributions will be calculated within each group to ensure that the distribution of the LLM agents is balanced; if the statistics are not balanced, the Testing Preparation Module will re-execute agent traffic allocation until satisfactory.
+These web environments for both groups need to be launched using independent browser instances controlled through ChromeDriver (for web environment parsing) and Selenium WebDriver (for automated interaction) integration.
 During the automatic interaction with the web environment, each LLM agent begins to interact with the webpage using an autonomous action prediction loop, which is shown in Fig.
 3.
-Each loop involves perceiving the current webpage state, interpreting the action space, predicting the next action, and executing that action in the browser.
-The details of this core automated agent-environment interaction process are explained below in Section 4.2.
+Each loop involves perceiving the webpage state, interpreting the action space, predicting the next action, and executing that action in the browser.
+.
 During the process, each step of the interaction is recorded, and the system monitors the overall session progression until termination, such that the agent accomplishes the intended goal or encounters failure cases.
@@
 A single experimental run can involve hundreds or thousands of sessions distributed across different personas and design conditions, all executed without any human intervention.
-Figure 3.
-One action prediction iteration of the automated web testing in Agent A/B with an LLM agent.
-(1) An Agent Profiling Module maintains a comprehensive agent description with an LLM-generated persona, user-specified intention, and the action history of the current session.
-In the meantime, (2) the Environment Parsing Module parses the webpage into structured web representation and action spaces.
-(3) All the information is fed into the LLM Agent for the next action prediction, (4) which will be executed by the Action Execution Module in the web environment to drive the next iteration.
-:
-4.2. Agent-Environment Interaction Architecture
+A.2. Agent-Environment Interaction Architecture
 At the core of Agent A/B is an iterative mechanism where each LLM agent continuously interacts with a real web environment by dynamically updating its understanding of the environment and adjusting its actions accordingly.
@@
 Together, these components enable robust operation in complex and dynamic web environments, for instance, our system evaluation in Section 5 is experimented on the Amazon [^1] http://www.Amazon.com platform.
+We will release our code upon acceptance of this paper.
 Environment Parsing Module.
@@
 On the search result page, as shown in Fig.
-4, for instance, we extract product details such as titles, names, ratings, reviews, and prices from the results section and gather filter options (e.g., Brand, Price, Delivery Day) from the left search panel.
+7, for instance, we extract product details such as titles, names, ratings, reviews, and prices from the results section and gather filter options (e.g., Brand, Price, Delivery Day) from the left search panel.
 This approach is designed to remove irrelevant elements that are not relevant to the targeted design features and user interactions, such as advertisements, banners, or unrelated hyperlinks.
@@
 In particular, the LLM agent models the next-step decision-making problem as a form of language-based reasoning and planning task by mapping structured state observations into reasoning traces and action predictions.
-Agent A/B is not bound to a specific LLM agent implementation.
+Agent A/B is not bound to a specific LLM agent.
 Instead, our system treats the LLM agent as an exchangeable module that supports various types of LLM web agents (ReAct (Yao et al., 2023), FireClaw [^2] https://github.com/mendableai/firecrawl-mcp-server) with convenient “plug-and-play” APIs, analogous to the Model Context Protocol (MCP) proposed by Claude.
-In our experiment, we adopt the UXAgent framework (Lu et al., 2025b), a state-of-the-art LLM agent for web interactions, as a representative implementation due to its strong performance and support for multi-step planning and intermediate memory.
-Figure 4.
-The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B.
-In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the users’ search query (e.g., “solar filter for telescope”).:
+In our experiment, we adopt the UXAgent framework (Yuxuan Lu et al., 2025), a state-of-the-art LLM agent for web interactions, as a representative implementation due to its strong performance and its built-in support for multi-step planning, hierarchical control, and intermediate memory management.
+UXAgent (Fig.
+4) follows a dual-loop architecture composed of a Fast Loop for rapid perception–action cycles and a Slow Loop for deeper reasoning.
+Within the fast loop, the perception module converts raw signals from the browser connector into natural-language observations, which are then appended to the memory stream.
+The planning module generates short-horizon plans conditioned on these observations, and the action module executes the action.
+The slow loop complements this by incorporating a wonder module that generates speculative or exploratory thoughts, simulating human-like mind wandering, and a reflection module that performs deliberate reasoning over past trajectories and current goals.
+Both loops continuously read from and write to the unified memory stream, allowing perception, planning, action, wonder, and reflection to jointly shape the agent’s evolving internal state and decision process.
+Figure 4. UXAgent Architecture design (Yuxuan Lu et al., 2025).
+The architecture features two interacting processes: the Fast Loop and the Slow Loop, which handle rapid web interactions and deeper reasoning respectively.
+Within the fast loop, the perception module converts raw observations from the environment into natural language memory entries stored in the memory stream, while the planning module generates short-term plans that the action module executes.
+In the slow loop, the wonder module simulates human “mind wandering” and the reflection module performs more thorough reasoning about the current context.
+Modules retrieve relevant entries from the memory stream and also append newly generated memories back into it.
 Action Execution Module.
@@
 Each completed session produces a full trace of interaction history, action rationales, page states, and final outcomes.
-5. Case Study: A/B Testing with Reduced Filter Options on Amazon.com
-To evaluate the effectiveness of Agent A/B, we conducted large-scale simulations on a live e-commerce platform.
-5.1. Method: A/B Testing Scenario
+Appendix B A/B Testing Scenario
 We designed a simulated A/B testing scenario focused on a novel feature design of the left-side filtering panel of Amazon’s shopping interface, a common area for UI experimentation in e-commerce.
-The goal was to evaluate how variations in filter configurations influence user behaviors such as dollar spending and product purchases.
-Control Condition
-Full Filter List
-Human, N=1M Control Condition
-Full Filter List
-Agent, N=500 Treatment Condition
-Reduced Filter List
-Agent, N=500
-Search 6.40 1.42 1.43
-Click_product 6.96 1.87 2.09
-Click_filter_option 0.33 0.58 0.60
-Purchase 0.62 0.81 0.83
-Stop - 0.19 0.17
-Average actions 15.96 6.05 6.60
-# of purchase * - 403 414
-Average \$ spend - \$55.14 \$60.99
-Table 2.
-Comparison of human customers’s actions per session in control condition (current design with full filter list), and the LLM agents as virtual customers in both control condition and treatment condition (new design with reduced filter list).
-The LLM agents in treatment condition purchased significantly more items than the ones in the control condition.
-$\chi^{2}(1)=5.51$, p-value ¡ 0.05
 Our team has the privilege of altering the left filter panel interface (Fig.
-4).
-In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the users’ search query.
-Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and randomly sampled 1,000 to simulate individual shopping sessions, 500 assigned into the treatment condition and the other 500 assigned into the control condition.
-In parallel, we also conducted the real human online A/B testing with 1 million users randomly assigned into the control condition and another 1 million users in the treatment condition.
+7).
+In comparison to the existing design as the control condition, where all the filter options are shown, the new design (treatment condition) uses a similarity-based ranking algorithm to reduce the filter options that have a lower-than-80% similarity score to the user’s search query.
+Using the LLM Agent Generation Module, we created 100,000 agent personas [^3] We will open source these agent personas and then randomly sampled 1,000 of them to simulate individual shopping sessions, with 500 assigned to the treatment condition and 500 to the control condition.
+The personas’ demographic attributes were sampled to match real online shopper distributions, including age (18–24: 15.8%, 25–34: 24.6%, 35–44: 22.1%, 45–54: 19.6%, 55–64: 17.9%), gender (49.4% male, 50.6% female), and household income (0–30,000: 32.1%, 30,001–94,000: 32.4%, 94,001–1,000,000: 35.5%).
+In parallel, we also conducted the real human A/B testing with 2 million users randomly assigned into the control condition and the treatment condition.
 Each agent was initialized with a persona profile and a shopping task (e.g., “find a budget smart speaker under \$40 with strong customer reviews”).
@@
 We logged full action traces, metadata (e.g., duration, outcome), and the agent’s rationales where applicable.
-5.2. Finding: Alignment with Human Behavior
-To benchmark our simulation system in aligning with human behavior, we conducted a simulated A/B testing case study using Agent A/B and also have access to the results of online A/B testing results with real human subjects using identical task conditions.
-Table 2 presents the aggregated actions and statistics from sessions with the 1,000 LLM agents participating and with 2M human participants [^4] Due to company policy, the human participant’s behavior data in the treatment condition can not be shared..
-As shown in Table 2, human participants in the control condition engaged in longer, more exploratory interactions than agents, averaging nearly twice as many actions per session (15.96 vs.
-6.05).
-They also clicked on more product detail pages (6.96 vs.
-1.87) and performed more searches (6.40 vs.
-1.42).
-Despite these differences in interaction style, humans and agents showed similar purchase rates (0.62 vs.
-0.81) and comparable use of filter options (0.33 vs.
-0.58).
-These findings highlight notable behavioral differences: while LLM agents adopt more goal-directed, structured interaction strategies, human users display broader exploratory behaviors.
-Nevertheless, the decision-making processes and general interaction patterns of the agents reasonably approximate human behavior for intention-driven tasks.
-This alignment supports the utility and validity of employing agent simulations for controlled UX evaluations, especially in scenarios where rapid interface assessment without live deployment is beneficial.
-Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.:
+To learn the environment cost of LLM Agents (as suggested by (Inie et al., 2025)), we calculated the number of tokens and CO2 emissions of our experiment.
+For 1,000 agent simulations using Claude 3.5 Sonnet, estimated 875 million tokens, resulting in a cost of approximately \$2,925 and CO2 emissions of 35 kilograms (using 0.00004 g CO2/token for average data center grid mix).
+In comparison, recruiting 1,000 human participants into a typical UX study session in the AI Industry would cost \$100 per person (\$100,000 in total).
+Figure 5. Number of LLM agents who completed a purchase under control and treatment conditions.
+A bar chart comparing the number of LLM agents who made a purchase.
+The control group with all filter options has 403 purchases, while the treatment group with reduced filter options has 414 purchases.
 (a) Average customer spending by gender groups across conditions.
 (b) Average customer spending by age groups across conditions.
-Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).:
-5.3. Findings: System Effectiveness Across Interface Variants
-To assess whether Agent A/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations.
-As shown in Table 2, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
-Agents in the treatment condition clicked on more products (2.09 vs.
-1.87), and the sessions have a higher total number of actions (6.60 vs.
-6.05, t(998)=1.08, p=0.28).
-Most notably, the LLM agents in the treatment condition click filter options more frequently than the ones in the control condition (0.60 vs.
-0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability.
-When looking at the final outcomes of the sessions in the treatment and control conditions, LLM-agent simulated virtual customers purchased more products in the treatment condition compared to the control condition(414 v.s.
-404, $\chi^{2}(1)=5.51$, p=0.03).
-This represents a modest but statistically significant increase in purchase completion, suggesting that reducing the number of filter options improved the likelihood of customers making a purchase.
-Expressed in practical terms, the treatment condition resulted in roughly a 2–3% higher conversion rate, which could translate into millions of additional purchases per day on Amazon.
-In terms of average spending across all customers, LLM agent simulated virtual customers in the treatment condition spent somewhat more than those in the control condition (\$60.99 v.s.
-\$55.14).
-However, this difference was not statistically significant (t(998) = 0.39, p = 0.69).
-The mean difference of about \$6 is small relative to the variability in spending, suggesting that while the treatment may have had a slight positive effect on purchase amounts, substantially larger sample sizes would be required to reliably detect such an effect.
-Breaking down average spending by gender, the data show different patterns across groups.
-Female customers spent an average of \$49.18 in the control condition and \$51.30 in the treatment condition, indicating only a small increase under the treatment.
-Male customers, by contrast, showed a larger increase, from \$61.82 (control) to \$70.83 (treatment).
-While statistical tests were not conducted for each subgroup separately, the direction of change suggests that the simplified filter interface may have a stronger positive effect on male customers’ spending than on female customers.
-A clearer set of differences emerges when comparing average spending across age groups.
-Among younger customers ($\leq 35$ ), average spending actually decreased in the treatment condition compared to the control condition ($\mathdollar 56.68\rightarrow\mathdollar 43.41$).
-Middle-aged customers (35–55) spent more under treatment ($\mathdollar 59.21\rightarrow\mathdollar 71.19$), while older customers (¿55) showed the largest increase ($\mathdollar 40.24\rightarrow\mathdollar 79.07$).
-These results suggest that the reduced filter interface may be particularly beneficial for older customers, potentially because it helps reduce choice overload.
-Conversely, younger customers may rely more on detailed filtering to find products that match their preferences, leading to lower spending when filter options are reduced.
-In sum, reducing the number of filter options was associated with a higher purchase completion rate and possible increases in customer spending, though the latter effects did not reach statistical significance.
-Subgroup analyses suggest heterogeneity in response: males and older customers appear to benefit more from the simplified interface, while younger customers may prefer more filtering flexibility.
-Given the variability in spending outcomes, further testing with larger samples will be important to confirm these trends.
-Notably, the LLM-agent A/B testing results aligned directionally with patterns observed in human A/B testing .
-This convergence increases confidence that agent-based simulations can anticipate real user behaviors and provide early insight into interface performance.
-Overall, the A/B simulation conducted with Agent A/B shows that LLM-based agents, when guided by personas and embedded intentions, generate behaviorally rich and interface-sensitive interactions.
-These behaviors vary in predictable ways across personas, align with humans, and detect subtle differences in design.
-Such findings support the system’s role as a behaviorally grounded, scalable complement to conventional A/B testing.
-The results suggest Agent A/B is well-suited for early-stage prototyping, pre-deployment validation, and hypothesis-driven UX research.
-6. Discussion
-Our design and evaluation of Agent A/B demonstrate the feasibility and utility of using LLM-based autonomous agents at scale for behavioral simulation in the life-cycle of the web interface designs.
-These findings echo a growing trend in Machine Learning and HCI, where large-scale LLM agent-based simulations are increasingly explored as tools for social simulations, economic experiments, and iterative design support (Park et al., 2023; Horton, 2023; Gao et al., 2024).
-In this section, we discuss the broader implications of our findings, system-level affordance, current limitations, and future research directions.
-6.1. Accelerating Design Iteration with Simulated Feedback
-One of the most immediate and impactful contributions of Agent A/B lies in its ability to expedite the feedback loop for A/B testing and design evaluation in web-based applications.
-Traditional A/B testing, despite remaining the gold standard for user-centered and evidence-based design validation approaches, suffers from well-documented challenges that are primarily related to the experiment participants.
-In particular, real user traffic is often limited and expensive, feedback is slow due to the long time span for experiment execution, and there are not enough resources to test all the design variants (Fabijan et al., 2017; Kohavi et al., 2013; Xu et al., 2015).
-These structural constraints make it difficult for teams to validate novel ideas, particularly during early prototyping stages, when feedback is most critical.
-Furthermore, many practical challenges in engineering time and testing infrastructure often restrict the number of A/B testings that can be executed.
-Recent work has emphasized the importance of early-stage, actionable feedback in early design phases to support interface and feature design (Fabijan et al., 2018).
-Tools like Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) have emphasized the value of rapid prototyping.
-However, most of these tools focus on interface prototyping rather than enabling feedback that is grounded in realistic user behavior.
-Our Agent A/B, on the contrary, fills this critical gap by introducing a new phase into the design life-cycle: agent-based pilot experiments before real user testing.
-Our system enables a rapid and automated feedback loop through two key technical capabilities: 1) automatically scaling up massive amounts of LLM agents with realistic and diverse personas, and 2) asking the agents to perform realistic behaviors by “role-playing” real humans.
-These simulations are executed within live browser environments, with LLM agents continuously perceiving structured page representations, reason through goals and interface states, and predicting plausible next actions.
-The predicted actions are automatically translated into executions in web environments to picture individual user journeys.
-As a result, Agent A/B users, such as UX designers, can obtain large-scale, fine-grained insights into user decision strategies with rationales, failure cases, and behavioral divergence across design variants.
-These insights can support end users in iterating the designs and revising the actual user A/B tests before any real human participants are recruited or any feature is deployed.
-6.2. Inclusive Piloting and Risk-Free Testing for Underrepresented Populations
-In addition to accelerating iteration, Agent A/B enables another critical capability: inclusive, risk-free piloting for user populations that are otherwise difficult to recruit or ethically sensitive to test.
-These populations (e.g., older adults) are often underrepresented in early testing phases due to ethical or logistical barriers, and a poorly designed application may even impose harm on the user participants.
-Agent A/B allows designers to simulate interactions from such groups by configuring agent personas that reflect the demographic and behavioral characteristics of targeted populations.
-For example, agents can be instantiated to simulate limited digital literacy, slower decision-making, or specific accessibility needs to reflect the behaviors of older adults more realistically and believably.
-By using our Agent A/B to generate a targeted agent population at scale and, more importantly, to automatically and iteratively test different design variants, designers can not only better evaluate how these groups might experience proposed features in real user testing but also fix design flaws and avoid any potential risks.
-This capability not only enhances safety and inclusion in the design process but also ensures that interface and feature designs are more robust and accessible.
-Our idea echoes the early exploration of scaling interface evaluation using crowd-based approximations of end-user behavior (Bernstein et al., 2011; Oppenlaender et al., 2020).
-Our findings suggest that even in early prototyping phases, agents can offer intent-driven behavior patterns that may disclose usability bottlenecks, task friction, or potential failures.
-This paradigm shift reframes LLM agent simulation not merely as a performance benchmark but as a core design methodology.
-Agent-based piloting complements traditional user testing with risk-free experimentation, broadened design coverage, and prompt usability feedback that aligns with realistic interaction patterns.
-6.3. Behavioral Fidelity of LLM Agents
-Our results show that LLM-based agents can generate realistic behavior trajectories that are aligned with various human shopping strategies.
-The alignment in task completion rates, refinement behaviors, and interaction flow with real human data reflects the growing consensus in HCI that advanced LLMs can perform believable and context-aware behaviors (Park et al., 2023; Horton, 2023; Lu et al., 2025a).
-However, simulated behavior is not a complete substitute for human cognition.
-As prior work in cognitive modeling and HCI simulation has emphasized (Gray and Altmann, 2001; John and Kieras, 1996; Ziebart et al., 2009), real humans are influenced by complicated explicit and implicit factors, including emotion, fatigue, prior experience, and latent goals.
-Many of these factors are not accessible to the current LLM agents.
-For instance, in our evaluations, agent behavior tended to be more deterministic, focused, and efficient compared to human users, who often displayed broader exploration and casual browsing.
-The behavior trajectories, as well as reasoning generated by LLM agents, should not be interpreted as perfect replications of user behaviors but as structured approximations that emphasize user behavior coverage and consistency.
-Our idea is analogous to the use of cognitive and computational models in prior HCI work, such as GOMS and ACT-R, which were not designed to replicate human behaviors perfectly but to estimate the interaction trade-offs at scale (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996).
-6.4. Toward Simulation-Supported Design Methodologies and Automated Design Optimization
-While Agent A/B currently focuses on simulating A/B testing between interface design variants, our system architecture supports broader capabilities that move beyond this particular evaluation methodology.
-Our system laid the foundation for AI-based human behavior simulation at scale to support various design explorations, refinements, and evaluations.
-For instance, we can adapt our system to cognitive walkthrough and heuristic evaluation by designing the LLM agents to role-play different types of stakeholders (e.g., domain experts, HCI designers, technical developers, etc.) and specifying the agent’s intention to be aligned with the design methodologies correspondingly.
-This novel opportunity echoes a number of established works that underscore the development and adoption of automated tools for interface exploration and optimization promptly, such as Fuse (Kuznetsov et al., 2022), Apparition (Lasecki et al., 2015), and d.tools (Hartmann et al., 2006).
-Further, our Agent A/B sheds light on the vision of “design mining” (Kumar et al., 2013), where large numbers of alternatives are evaluated computationally to identify the most promising solutions.
-Agent A/B supports this approach by serving as a simulation engine for a large pool of candidate variants and automating the design optimization through full interaction traces.
-For example, if $50$ layout variants of a filter panel were proposed, agent simulations could conduct large-scale simulations and rank the designs by search efficiency, user engagement, or failure rate across different user subgroups.
-This extension of the system functions beyond as a testbed but as an interactive co-design partner to provide empirical evidence throughout the life-cycle of novel designs and functionalities.
-We view this direction as a natural evolution of the current system architecture and an exciting opportunity for future integration with generative design tools and adaptive interface frameworks.
-6.5. Limitations and Future Directions
-While Agent A/B establishes a strong foundation for the scalable and automated LLM agent-based A/B testing, we acknowledge two key limitations in the current scope of the system and inform ongoing development.
-First, the robustness of our system and the behavioral fidelity of simulated agents remain bound by the reasoning and grounding capabilities of the underlying LLMs.
-Although recent LLMs such as Claude 3 and GPT-4o exhibit exceptional language understanding and planning capacity, they can still misinterpret complex or unconventional DOM structures, particularly in dynamically rendered web environments.
-Even state-of-the-art LLMs may fail to correctly parse web environments when encountering inconsistent element labeling, interactive latency, or unexpected modal interruptions (Zhou et al., 2023; Koh et al., 2024).
-While our system mitigates these issues through a novel implementation of leveraging JSON for structured environment parsing, LLM agent performance can still degrade when real-time content, along with the web environment structures, continuously updates and changes.
-Second, agent behaviors in the current Agent A/B system are not able to comprehensively incorporate affective or meta-cognitive signals such as uncertainty, fatigue, or emotional response.
-These dimensions could be challenging for LLMs to model but are critical in shaping human cognitive behaviors and decision-making (Horton, 2023; Ziebart et al., 2009).
-LLM agents in our system tend to simulate goal-driven behavior with plausible cognitive structure but do not yet capture the full range of human variability or intent ambiguity.
-Beyond these limitations, several technical and conceptual opportunities emerge for future work.
-One promising direction lies in expanding agent capabilities through multimodal information perception.
-For instance, integrating visual inputs (e.g., screenshots, spatial layout) with text-based content could allow agents to operate more robustly across richly designed or accessibility-diverse interfaces but still may introduce irrelevant information along the path.
-Recent progress in vision-language models and multimodal agents (Gupta and Kembhavi, 2023; Chezelles et al., 2024) suggests a pathway to generalizing agent behaviors beyond text-only interfaces.
-Another important opportunity is collaborative or multi-agent simulation, especially for applications such as shared productivity tools, learning platforms, or social platforms.
-In these scenarios, agent interactions may involve negotiation, coordination, or conflict.
-Building on work like generative agents for social simulations (Park et al., 2023), future systems could simulate rich user ecosystems rather than independent task-solving to better reflect the collaborative work nature in real-world scenarios.
-These directions demonstrate the extensibility of Agent A/B and its potential to evolve from a simulation platform into a broader foundation for behaviorally informed, user-model-driven design tools.
-As LLM-based interaction models continue to improve, we anticipate these agents will not only validate designs but also help co-create them.
-7. Conclusion
-In this paper, we presented Agent A/B, a system that enables large-scale, LLM agent-based simulation of A/B testing for web interfaces in real browser-based web environments.
-Our evaluation demonstrated that LLM agents exhibit realistic, goal-aligned behaviors, are sensitive to interface variations, and provide actionable feedback comparable to human users.
-By supporting rapid, risk-free behavioral piloting, Agent A/B introduces a new phase for agent-based piloting in the design life-cycle that complements traditional A/B testing and expands the scope of early-stage UX evaluation.
-We envision future extensions that further enhance agent fidelity, broaden domain coverage, and integrate simulation into intelligent design optimization workflows.
-Appendix A Example of Persona Generated by Agent A/B
+Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).
+[…]
+Participant Gender Job Title Team
+P1 F Project Manager PM Team
+P2 M Software Development Manager Engineer Team
+P3 F Product Manager PM Team
+P4 M Product Manager PM Team
+P5 M Product Manager PM Team
+P6 M Machine Learning Researcher Science Team
+Table 2. Demographics of formative study participants.
+Appendix C Example of Persona Generated by Agent A/B
 Persona: Marcus
-[…]
+Background:
 Marcus is a 35-year-old freelance graphic designer living in Austin, Texas.
 After working for a decade in various creative agencies, he transitioned to freelancing to gain more control over his schedule and focus on passion projects, such as illustrating indie game assets and creating digital art for local musicians.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Demographics:
+Age: 35
+Gender: Male
+Education: Bachelors degree in Visual Communication
+Profession: Freelance Graphic Designer
+Income: \$70,000 (variable based on projects)
+Financial Situation:
 Marcus earns a decent living from his freelance gigs, though his income can fluctuate.
@@
 He sets aside part of his earnings for travel and software upgrades, which are essential for his work.
-[…]
+Shopping Habits:
 Marcus enjoys discovering unique or niche products, especially tech gadgets, art supplies, and streetwear.
@@
 For clothing, hes drawn to bold, graphic-heavy items that reflect his artistic vibe.
-[…]
+Professional Life:
 Marcus works from a home studio that doubles as a creative space.
@@
 He frequently updates his portfolio and maintains a strong social media presence to attract new clients.
-[…]
+Personal Style:
 Marcus has an edgy and expressive fashion sense.
@@
 His go-to outfit is a soft hoodie with a custom design, black joggers, and high-top sneakers.
+Appendix D Persona Generation Prompt
+You are a helpful assistant that generates diverse personas.
+Examples:
+[PERSONA EXAMPLE]
+Generate a persona using the above examples. The persona should be different from previous personas to ensure diversity. The persona should:
+- Have the age of {age}
+- Be {gender}
+- Have an income between \${income_range[0]} and \${income_range[1]}
+Provide the persona in the same format as the examples.
+Only output the persona, no other text.
+Figure 7.
+The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B.
+In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the user’s search query (e.g., “solar filter for telescope”).
+A side-by-side comparison of two Amazon search results pages for the query ”solar filter for telescope.” Both pages show similar products but have differences in their filter options on the left sidebar.
+The left image displays additional filter categories under ”Department” such as ”Telescope Refractors,” ”Safety Goggles & Glasses,” and ”Telescope Eyepieces,” which are not present in the right image.
+Both images include nested subcategories under ”Camera & Photo Filters” and identical product listings on the right side of the page.
+This comparison highlights subtle differences in the user interface that could affect user navigation or filtering behavior.
+Appendix E Interview Protocol
+1. Overall Experimentation Practice
+Can you describe your general workflow for running A/B testing experiments in your company?
+2. End-to-End Workflow and Timing
+How does an experiment progress from an initial design idea to a final deployment decision, and how long do the major stages typically take?
+3. Pain Points and Challenges
+What are the main challenges or friction points you encounter when designing, running experiments or analyzing results?
+4. Experimentation Tools and Platform Evaluation
+How well does the existing A/B testing experiment tool support your work, and what new features would you want?
+5. Other
+Is there anything else you would like to add, or do you have any questions for me?